### 03 May 2026 v1

# Separability from Multipartite Measures

Separability from Multipartite Measures
Chen-Te Maa,111e-mail address: yefgst@gmail.com and Ma-Ke Yuanb,222e-mail address: mkyuan19@fudan.edu.cn
a Department of Physics, Great Bay University, Dongguan, Guangdong 52300, China.
b Department of Physics and Center for Field Theory and Particle Physics,
Fudan University, Shanghai 200433, China.
We show that the third-order negativity provides a necessary and sufficient criterion for full separability of tripartite pure states, and extend this to mixed states beyond bipartite diagnostics such as negativity. As a minimal nontrivial example, a four-qubit pure state has three-qubit mixed reductions; its complete characterization requires six bipartite, eight tripartite, and four quadripartite measures, with the third-order negativity serving as a key separability criterion. We further generalize these separability criteria to multipartite qudit systems and discuss an application to conformal field theory.
1 Introduction
Quantum entanglement is a phenomenon in which the quantum state of a composite system cannot, in general, be described independently in terms of its subsystems, even when large distances separate them. In quantum information theory, deciding whether a density matrix is separable, namely, whether it can be written as
$\rho=\sum_{j}p_{j}\rho_{1}^{(j)}\otimes\rho_{2}^{(j)}\otimes\cdots\rho_{n}^{(j)}\ ,$ (1)
is known as the separability problem. Although separable states should contain correlations, these correlations are purely classical rather than quantum. In general, the separability problem is believed to be computationally hard.
A useful strategy is therefore to construct entanglement measures whose vanishing gives a separability criterion. The simplest example is the bipartite two-qubit pure state, for which the two-tangle
$\tau_{A|B}=2\big(1-\operatorname{Tr}(\rho_{A}^{2})\big)$ (2)
vanishes if and only if the state is a product state [3, 24, 49]. Its key ingredient is the reduced purity $\operatorname{Tr}(\rho_{A}^{2})$, which reaches its maximal value $1$ precisely for product states and is symmetric under exchanging $A$ and $B$. The convex-roof extension,
$E_{\tau_{A|B}}(\rho)\equiv\inf_{\{p_{\ell},|{\psi_{\ell}}\rangle\}\in\mathbb{D(\rho)}}\sum_{\ell}p_{\ell}\,\tau_{A|B}(\psi_{\ell})\ ,$ (3)
with $\mathbb{D}(\rho)$ the set of all convex decompositions of $\rho$ into pure states, then provides a natural mixed-state separability criterion [3, 24, 49]. This motivates the search for an analogous tripartite quantity: a fully permutation-symmetric measure that vanishes precisely on fully product pure states, admits a natural convex-roof extension to mixed states, and can therefore serve as a useful criterion for full separability.
Three-qubit pure states provide the simplest nontrivial setting for genuine tripartite entanglement [12], already exhibiting features absent in the bipartite case. In two-qubit systems, any pure state can be brought by local unitary (LU) transformations $\mathrm{SU}(2)^{\otimes\mathtt{2}}$ to the Schmidt form
$|\psi\rangle=\lambda_{0}|00\rangle+\lambda_{1}|11\rangle\ ,\quad\text{with}\ \ \lambda_{j}\geq 0\ ,\ \lambda_{0}^{2}+\lambda_{1}^{2}=1\ ,$ (4)
reflecting the fact that a single parameter completely characterizes its bipartite entanglement. However, such a decomposition does not generalize straightforwardly to multipartite systems. For three-qubit systems, using the generalized Schmidt decomposition [34], any pure state is LU-equivalent to the canonical form [1, 4]
$|\psi\rangle=\lambda_{0}|000\rangle+\lambda_{1}e^{i\varphi}|100\rangle+\lambda_{2}|101\rangle+\lambda_{3}|110\rangle+\lambda_{4}|111\rangle\ ,$ (5)
with $\lambda_{j}\geq 0$, $0\leq\varphi<\pi$, and $\sum_{j=0}^{4}\lambda_{j}^{2}=1$. This already shows the increased structural complexity of tripartite entanglement. Beyond bipartite measures, two important genuinely tripartite quantities arise [37]:
  • Three-Tangle [12]
    $\tau_{ABC}=\tau_{A|BC}-\tau_{A|B}-\tau_{A|C}\ ,$ (6)
    which captures the residual entanglement that cannot be reduced to pairwise correlations. Here $\tau_{A|BC}=2(1-\operatorname{Tr}(\rho_{A}^{2}))$, while $\tau_{A|B}$ and $\tau_{A|C}$ are defined via the convex-roof extension for mixed states, i.e., Eq. (3).
  • Third-Order Negativity [37]
    $I_{5}=\operatorname{Tr}\!\big[(\rho_{BC}^{\Gamma})^{3}\big]\ ,\quad\rho_{BC}=\operatorname{Tr}_{A}|\Psi\rangle\langle\Psi|\ ,$ (7)
    where $\Gamma$ denotes the partial transpose. Although $\rho_{BC}^{\Gamma}$ is not positive in general, it is Hermitian, so $I_{5}$ is always real.
For three-qubit pure states, $I_{5}$ is fully symmetric under permutations of $A$, $B$, and $C$, making it a natural candidate for a permutation-invariant entanglement diagnostic [37]. Therefore, $\tau_{ABC}$ and $I_{5}$ illustrate that multipartite entanglement contains structures that cannot be captured solely by bipartite measures.
The Positive Partial Transposition (PPT) criterion provides a powerful tool for detecting quantum entanglement in bipartite systems [35]. For a density matrix $\rho$, one considers its partial transpose. If $\rho$ is separable, then all eigenvalues of its partial transpose are nonnegative [35]. Therefore, the presence of any negative eigenvalue signals quantum entanglement [35]. This idea underlies negativity [44] and logarithmic negativity [36], which are widely used entanglement measures [42, 45]:
  • They vanish on separable states;
  • They are monotonic under the local operations and classical communication (LOCC) [31]
    $E\Big(\sum_{j}p_{j}\rho_{j}\Big)\leq E(\rho)\ .$ (8)
Moreover, negativity and logarithmic negativity have several practical advantages: they are computable for mixed states, both analytically and numerically [11], without requiring a convex-roof construction [36], and they are relatively accessible in experiments [41]. These features make it especially useful to study the Rényi negativities [13, 50], whose spectra encode richer information about entanglement. However, the PPT criterion has an important limitation: It is necessary and sufficient only in low dimensions, specifically for $2\times 2$ and $2\times 3$ Hilbert spaces [26]. In higher dimensions, there exist entangled states that are PPT but still entangled [26]. Thus, PPT is no longer a sufficient criterion for separability in general.
In multipartite systems, the situation becomes even more challenging: determining separability typically requires examining all possible bipartitions, which is highly cumbersome both theoretically and experimentally. This motivates the search for multipartite entanglement measures that:
  • provide a single, unified criterion for separability,
  • are permutation symmetric,
  • admit a computable formulation, and
  • ideally correspond to a single observable or measurement protocol.
Establishing such measures—such as the third-order negativity $I_{5}$—offers a promising route to simplifying the detection of multipartite entanglement and to understanding its structure beyond bipartite correlations.
The third-order negativity $I_{5}$ has two notable advantages: first, it is experimentally accessible, for example via randomized measurements or tomography-inspired protocols [9, 16]; second, for three-qubit pure states, saturation of its lower bound identifies the W-state [33], providing a clear physical interpretation and a way to distinguish entanglement classes. However, a key drawback is that $I_{5}$ does not satisfy the full set of requirements for a proper entanglement measure and is therefore not an entanglement measure in its own right. To address this issue, Ref. [32] introduced the quantity
$\phi_{ABC}=69-\operatorname{Tr}\big[(2\rho_{AB}+\rho_{A}\otimes I_{B}+I_{A}\otimes\rho_{B})^{3}\big]-3\operatorname{Tr}(\rho_{AB}^{2})\ ,$ (9)
which is an entanglement measure. Using identities specific to three-qubit pure states, we show that this quantity can be rewritten in terms of $I_{5}$ and the tangles:
$\phi_{ABC}=12(1-I_{5})+27(\tau_{A|B}+\tau_{A|C}+\tau_{B|C})+\frac{81}{2}\tau_{ABC}\ .$ (10)
Eq. (10) reveals several key features:
  • Permutation symmetry. The expression (10) is manifestly symmetric under $A$, $B$, $C$, which is essential for a genuine multipartite measure.
  • Decomposition of entanglement structure. The measure separates into:
    • $I_{5}$: a global, higher-order correlation (experimentally accessible),
    • pairwise tangles: bipartite entanglement content,
    • $\tau_{ABC}$: genuine tripartite entanglement.
  • Physical interpretation. $\phi_{ABC}$ effectively combines:
    • detectability (via $I_{5}$),
    • monotonicity (through its construction),
    • completeness (capturing both bipartite and tripartite structures).
Eq. (10) shows that although $I_{5}$ alone is not sufficient, it becomes part of a well-defined entanglement measure when combined with lower-order measures. In this sense, $\phi_{ABC}$ provides a bridge between experimentally measurable quantities and theoretically well-defined entanglement measures. More importantly, it supports our broader goal of constructing permutation-symmetric quantities that can diagnose multipartite separability and entanglement, which we now discuss in the general multipartite setting.
[…]
$E\Big(\sum_{j}p_{j}\rho_{j}\Big)\leq\sum_{j}p_{j}E(\rho_{j})\leq E(\rho)\ ,$ (11)
so that they remain valid entanglement measures under convex-roof extension [15]. This guarantees that our pure-state classification extends consistency to mixed states.
To figure out the 4 quadripartite measures, we first note that the four-tangle [40, 48]
$4|H|^{2}=C^{2}_{1234}(\Psi)\ ,\quad\text{with}\ \ C_{1234}(\Psi)=|\langle\Psi^{*}|\sigma_{y}^{\otimes 4}|\Psi\rangle|$ (12)
plays the role of a genuine quadripartite invariant, analogous to the two-qubit concurrence $\tau_{A|B}=C^{2}(\Psi)$ with $C(\Psi)=|\langle\Psi|\sigma_{y}^{\otimes 2}|\Psi^{*}\rangle|$. The key structural identity [21]
$4|H|^{2}=\tau_{A|BCD}+\tau_{B|ACD}+\tau_{C|ABD}+\tau_{D|ABC}-\tau_{AB|CD}-\tau_{AC|BD}-\tau_{AD|BC}$ (13)
shows that $4|H|^{2}$ is not determined solely by lower-order bipartite and tripartite correlations, but instead provides an independent quadripartite contribution. It is therefore a natural ingredient needed to complete the eighteen-parameter set of entanglement measures. For a positive homogeneous function $\mu$,
$\mu(\lambda\psi)=\lambda^{\eta}\mu(\psi)\ ,$ (14)
the condition that the degree $\eta\leq 4$ ensures that $\mu$ is an entanglement measure [17]. Since $4|H|^{2}$ is degree-4, it is an entanglement measure. Another useful observation is that both the two-tangle and the three-tangle can be rewritten in terms of a hyperdeterminant. Therefore, another natural candidate is the four-qubit hyperdeterminant $\Delta$, which is degree-24. Even though it is complicated, the key insight is: It can be expressed in terms of lower-degree invariants [38, 27, 30], including the degree-4 four-tangle $4|H|^{2}$, a degree-8 quantity $\Sigma$, a degree-12 quantity $\Pi$, and a degree-6 quantity $W$. We therefore identify $4|H|^{2}$, $|\Sigma|^{1/2}$, $|\Pi|^{1/3}$, and $|\Delta|^{1/6}$ as four independent quadripartite measures. See Table 1 for a summary.
The difficulty of classifying multipartite entanglement comes from the fact that the number of inequivalent entanglement classes grows combinatorially with qubit number. This complexity already appears in the simplest nontrivial cases: three qubits admit 2 inequivalent genuine entanglement classes [14], while four qubits admit 9 inequivalent families [43]. Thus, unlike bipartite systems, where quantities such as the entanglement entropy provide a complete ordering for pure states, multipartite entanglement cannot be totally ordered or fully quantified by a single scalar, but by a hierarchy of structures.
Different entanglement measures probe different aspects of correlations. For example, the three-tangle detects GHZ-type tripartite entanglement, $I_{5}$ is sensitive to W-type structures, and pairwise tangles capture bipartite correlations. Thus, two states may both be entangled while still being incomparable under a single scalar measure. This is one of the reasons why multipartite systems exhibit inequivalent entanglement classes. Because brute-force classification is impossible, the field focuses on representative families of states. A paradigmatic example is the generalized GHZ-state
$|\widetilde{\text{GHZ}}\rangle=a|00\cdots 0\rangle+b|11\cdots 1\rangle\ ,$ (15)
which exhibits purely global entanglement: after tracing out one or more qubits, all reduced density matrices become separable, and no local entanglement survives [39]. More remarkably, the converse direction also holds in an appropriate sense: detecting the separability of all reduced subsystems effectively identifies a GHZ-type structure. In local quantum field theory (QFT), the situation is even more extreme. The theory has infinitely many degrees of freedom, and its entanglement structure is further complicated by ultraviolet (UV) divergences that require regularization. Moreover, entanglement is generically present in spatial subregions [47]. Therefore, in local QFT, separability in position space is essentially lost. While position-space separability is lost in local QFT, it can re-emerge in momentum space. For example, the vacuum state in Lifshitz theory can exhibit mode-wise separability [2].
To make progress, one usually combines the replica trick [25], twist operators [6], and conformal field theory (CFT) [18, 28]. This framework provides a systematic way to compute Rényi entropy and Rényi negativity [5]. Conformal symmetry fixes the universal part of these quantities. At the same time, theory-dependent data, such as operator product expansion (OPE) coefficients [18, 28], can be obtained from anti-de Sitter/conformal field theory (AdS/CFT) correspondence [29, 22, 46, 10].
The AdS/CFT correspondence is a conjectured relationship between two distinct classes of physical theories. On one side lies Anti-de Sitter (AdS) space, which arises in theories of quantum gravity. In contrast, on the other side lie conformal field theories (CFTs), which are QFTs defined on the asymptotic boundary of spacetime. This idea is deeply motivated by insights from black hole thermodynamics [23], particularly the observation that the number of degrees of freedom associated with a region containing a black hole scales with its horizon area rather than its volume. This insight led to the broader holographic principle and, ultimately, to the AdS/CFT correspondence.
Although a complete theory of quantum gravity remains elusive, the AdS/CFT framework provides a powerful tool for studying strongly coupled quantum systems. In particular, it enables the exploration of semiclassical gravity and the large central-charge CFTs via the holographic dictionary [46]. In many cases, computations on the gravitational (AdS) side are technically simpler than direct calculations in the CFT, allowing one to probe regimes of QFT that are otherwise analytically inaccessible. The intrinsic difficulty of CFT and more general QFTs largely stems from the fact that they are defined on infinite-dimensional Hilbert spaces. The AdS/CFT correspondence offers a way to reorganize this complexity into a geometric framework, thereby providing new insights into strongly correlated quantum systems.
We can also use qudit systems as a controlled bridge to QFT. This works because the qudit systems mimic QFT subregions from the large local dimension. Therefore, they provide a clean laboratory for multipartite structures [19, 20] and separability criteria.
1.1 Summary of Results
This paper provides a comprehensive exploration of the relationship between separability and multipartite entanglement measures. We introduce the notion of third-order negativity and its multipartite generalizations, together with several applications. In this sense, our work establishes a concrete bridge between separability criteria and multipartite entanglement measures. Specifically:
2-qubit pure state 3-qubit pure state 4-qubit pure state
Bipartite $\tau_{A|B}$ $\tau_{A|B},\,\tau_{A|C},\,\tau_{B|C}$
$\tau_{A|B},\,\tau_{A|C},\,\tau_{A|D}$
$\tau_{B|C},\,\tau_{B|D},\,\tau_{C|D}$
Tripartite $\tau_{ABC},\,\phi_{ABC}$
$\tau_{ABC},\,\tau_{ABD},\,\tau_{ACD},\,\tau_{BCD}$
$\phi_{ABC},\,\phi_{ABD},\,\phi_{ACD},\,\phi_{BCD}$
Quadripartite $4|H|^{2},\,|\Sigma|^{1/2},\,|\Pi|^{1/3},\,|\Delta|^{1/6}$
Table 1: Hierarchical organization of entanglement measures for pure states of two, three, and four qubits.
  • We show that the third-order negativity $I_{5}$ attains its maximal value $1$ if and only if the state is fully product [7]. Motivated by the convex-roof construction in the bipartite case, we develop a corresponding mixed-state extension and obtain a necessary and sufficient criterion for full separability of tripartite mixed states. In this sense, $\phi_{ABC}$ (9) provides a natural tripartite analog of the two-tangle (2), while $I_{5}$ plays a role similar to the reduced purity $\operatorname{Tr}(\rho_{A}^{2})$ in Eq. (2). We also prove that, for three-qubit pure states, the saturation of the lower bound of $I_{5}$ and the upper bound of $\phi_{ABC}$ are achieved uniquely by the W- and GHZ-state, respectively.
  • The four-qubit pure state is the simplest nontrivial system whose one-qubit reductions yield tripartite mixed states, making it the natural arena for systematically applying the separability criterion. We identify a complete set of eighteen independent entanglement measures for four-qubit pure states, summarized in Table 1. Among them, $\phi_{ABC}$ and its permutations play a distinguished role: they are closely related to $I_{5}$ and provide necessary and sufficient criteria for full separability of tripartite states. Furthermore, we show that the remaining quadripartite measures— $4|H|^{2}$, $|\Sigma|^{1/2}$, and $|\Pi|^{1/3}$—are all generated by the hyperdeterminant $\Delta$, motivating the search for new entanglement measures derived from $\Delta$.
  • We demonstrate that among the eighteen measures, the four-tangle $4|H|^{2}$ is the only one that can remain nonzero when all the other seventeen measures vanish. This situation is realized uniquely by the generalized GHZ-type four-qubit states. Consequently, the entanglement of generalized GHZ-states can be quantified unambiguously.
  • We generalize the relation between separability and third-order negativity to generic finite-dimensional multipartite systems. We also show that the condition that all $(\mathtt{q}-1)$-partite reduced density matrices are fully separable is equivalent to the statement that the $\mathtt{q}$-partite pure state is LU-equivalent to a generalized GHZ-state. This provides a direct application of multipartite separability criteria to characterize the global entanglement structure. This proof also implies that the GHZ-type state is not allowable in the local QFT for the spatial bipartition.
  • Finally, we demonstrate an application in QFT by computing the third-order negativity in two-dimensional conformal field theory (CFT2) at large central charge, using the AdS/CFT correspondence to extract the relevant OPE coefficients.
The paper is organized as follows. In Sec. 2, we establish the relation between third-order negativity and tripartite separability. In Sec. 3, we introduce the complete set of eighteen entanglement measures and analyze their structural properties. In Sec. 4, we show that the four-tangle is the unique independently surviving measure. In Sec. 5, we generalize our results to multipartite systems. In Sec. 6, we present the CFT application of third-order negativity. Finally, Sec. 7 summarizes our results and discusses future directions. The Appendices A and B provide the detailed proof of the upper bound of $I_{5}$ and the lower bound of $\phi_{ABC}$.
2 Separability of Tripartite Systems
In this section, we show that the third-order negativity provides a criterion for full separability of tripartite systems. We begin with pure states, where the result takes a simple form: for a finite-dimensional tripartite pure state, $I_{5}=1$ if and only if the state is fully product. We then turn to three-qubit mixed states and obtain the criterion that $E_{\phi_{ABC}}=0$ if and only if the state is fully separable, where $\phi_{ABC}$ is a quantity closely related to $I_{5}$. Finally, based on $I_{5}$, we develop another mixed-state extension $E_{1-I_{5}}$, whose vanishing serves as a criterion for full separability in general finite-dimensional tripartite mixed states. We also show in Appendices A and B that for three-qubit pure states, the lower bound of $I_{5}$ is $2/9$, achieved by the W-state, and the upper bound of $\phi_{ABC}$ is $99/2$, achieved by the GHZ-state. Hence, one can determine how close a three-qubit pure state is to the W- and GHZ-state via the measurement of $I_{5}$ and $\phi_{ABC}$.
2.1 Tripartite Pure State
For tripartite pure states $|\Psi\rangle_{ABC}$, the third-order negativity $I_{5}$ (7) can also be expressed through the replica trick as
$I_{5}=\langle\Psi^{\otimes 3}|\Omega_{A}\Omega_{B}\Omega_{C}|\Psi^{\otimes 3}\rangle\ ,\ \ \text{with}\ \ \Omega_{A}=(1)(2)(3)\ ,\ \Omega_{B}=(123)\ ,\ \Omega_{C}=(132)\ .$ (16)
Here $\Omega_{A}$, $\Omega_{B}$, and $\Omega_{C}$ are permutation operators acting on the replica copies of the corresponding subsystems. In particular, $\Omega_{A}=(1)(2)(3)$ is the identity permutation on the three replica copies of subsystem $A$, while $\Omega_{B}=(123)$ cyclically permutes the three replicas of subsystem $B$, i.e.,
$\Omega_{B}(1)=2\ ,\quad\Omega_{B}(2)=3\ ,\quad\Omega_{B}(3)=1\ ,$ (17)
whereas $\Omega_{C}$ acts in the opposite direction. From Eq. (16) one can see that $I_{5}$ is symmetric among $A$, $B$, and $C$. We now prove the following proposition:
Proposition 1.
Let $|\Psi\rangle\in\mathcal{H}_{A}\otimes\mathcal{H}_{B}\otimes\mathcal{H}_{C}$ be a normalized 3-qudit pure state, then
$I_{5}=1\iff|\Psi\rangle\text{ is a fully product state}\ .$ (18)
Proof.
The converse is immediate. For the nontrivial direction, we first prove that for a general bipartite mixed state $\rho_{BC}$,
$I_{5}=\operatorname{Tr}\big[(\rho_{BC}^{\Gamma})^{3}\big]\leq 1\ .$ (19)
Let $X=\rho_{BC}^{\Gamma}$. Since $X$ is Hermitian, its eigenvalues $\mu_{j}$ are real, and
$I_{5}=\sum_{j}\mu_{j}^{3}\ .$ (20)
We now show that $\mu_{j}\leq 1$ for all $j$. It is clear that
[…]
Now let $|v\rangle=\sum_{k}s_{k}|k\rangle_{B}|k\rangle_{C}$ be the Schmidt decomposition of $|v\rangle$. Then the eigenvalues of $(|v\rangle\langle v|)^{\Gamma}$ are $s_{k}^{2}$ and $\pm s_{k}s_{l}$ with $k<l$. Hence, we obtain
$\lambda_{\max}\big((|v\rangle\langle v|)^{\Gamma}\big)=\max_{k}s_{k}^{2}\leq 1\Longrightarrow\lambda_{\max}(X)\leq 1\ ,$ (22)
so every eigenvalue $\mu_{j}$ of $X$ satisfies $\mu_{j}\leq 1$. It follows that $\mu_{j}^{3}\leq\mu_{j}^{2}$ for every $j$, and therefore
$I_{5}=\sum_{j}\mu_{j}^{3}\leq\sum_{j}\mu_{j}^{2}=\operatorname{Tr}(X^{2})=\operatorname{Tr}(\rho_{BC}^{2})\leq 1\ ,$ (23)
where the third equal sign follows from the fact that the partial transpose preserves the Hilbert-Schmidt norm. Now suppose that $I_{5}=1$. Since
$I_{5}\leq\operatorname{Tr}(\rho_{BC}^{2})\leq 1\ ,$ (24)
both inequalities must be saturated. In particular, $\operatorname{Tr}(\rho_{BC}^{2})=1$ implies that $\rho_{BC}$ is pure. Hence, $|\Psi\rangle$ factorizes across $A|BC$,
$|\Psi\rangle=|A\rangle\otimes|\phi\rangle_{BC}\ .$ (25)
For a bipartite pure state $|\phi\rangle_{BC}$, one has
$\operatorname{Tr}\!\big[\big((|\phi\rangle\langle\phi|)^{\Gamma}\big)^{3}\big]=\operatorname{Tr}(\rho_{B}^{3})\ ,\quad\rho_{B}=\operatorname{Tr}_{C}|\phi\rangle\langle\phi|\ .$ (26)
Therefore, we have $1=I_{5}=\operatorname{Tr}(\rho_{B}^{3})$, which implies that $\rho_{B}$ is also pure. Thus $|\phi\rangle_{BC}$ is product across $B|C$, and $|\Psi\rangle$ is fully product across $A|B|C$. ∎
Lower bound of $I_{5}$ for three-qubit pure state.
The lower bound is $I_{5}\geq 2/9$, which saturates when the state is W-state. See Appendix A for a proof [^3] Ref. [33] stated that the minimum of $I_{5}$ is attained by the W-state, with the value $2/9$. We have not found a proof of this statement in the existing literature.. Therefore, by measuring $I_{5}$, one can determine how close a 3-qubit pure state is to the W-state.
2.2 Three-Qubit Mixed State
We first give a derivation of Eq. (10), which rewrites the quantity $\phi_{ABC}$ (9) in terms of $I_{5}$. It has been shown in Ref. [32] that Eq. (9) can be expanded as
$\phi_{ABC}=69-12I_{5}-16\big[\!\operatorname{Tr}(\rho_{A}^{3})+\operatorname{Tr}(\rho_{B}^{3})+\operatorname{Tr}(\rho_{C}^{3})\big]-3\operatorname{Tr}(\rho_{A}^{2})-3\operatorname{Tr}(\rho_{B}^{2})-3\operatorname{Tr}(\rho_{AB}^{2})\ .$ (27)
For three-qubit pure states, $\operatorname{Tr}(\rho_{AB}^{2})=\operatorname{Tr}(\rho_{C}^{2})$. For normalized density matrix $\rho$ with rank no more than 2, we have $\operatorname{Tr}(\rho^{3})=(3\operatorname{Tr}(\rho^{2})-1)/2$. Therefore,
$\phi_{ABC}=93-12I_{5}-27\big[\!\operatorname{Tr}(\rho_{A}^{2})+\operatorname{Tr}(\rho_{B}^{2})+\operatorname{Tr}(\rho_{C}^{2})\big]\ .$ (28)
Using the definition of the two-tangle of pure states and Eq. (6), we have
$\begin{split}\operatorname{Tr}(\rho_{A}^{2})+\operatorname{Tr}(\rho_{B}^{2})+\operatorname{Tr}(\rho_{C}^{2})&=3-\frac{1}{2}(\tau_{A|BC}+\tau_{B|CA}+\tau_{C|AB})\\ &=3-(\tau_{A|B}+\tau_{A|C}+\tau_{B|C})-\frac{3}{2}\tau_{ABC}\ .\end{split}$ (29)
Substituting Eqs. (28) and (29) into Eq. (27), we obtain Eq. (10).
We now show that $\phi_{ABC}$ can serve as a necessary and sufficient criterion for full separability of three-qubit mixed states.
Proposition 2.
Let
$E_{\phi_{ABC}}(\rho)\equiv\inf_{\{p_{\ell},|{\psi_{\ell}}\rangle\}\in\mathbb{D}(\rho)}\sum_{\ell}p_{\ell}\,\phi_{ABC}(\psi_{\ell})\ ,$ (30)
where $\phi_{ABC}$ is defined as Eq. (10):
$\phi_{ABC}=12(1-I_{5})+27(\tau_{A|B}+\tau_{A|C}+\tau_{B|C})+\frac{81}{2}\tau_{ABC}\ .$
Then, for any three-qubit mixed state $\rho$,
$E_{\phi_{ABC}}(\rho)=0\iff\rho\ \text{is fully separable}\ .$ (31)
Proof.
One the one hand, since $\tau_{A|B}$, $\tau_{B|C}$, $\tau_{C|A}$, $\tau_{ABC}$ and $1-I_{5}$ are all nonnegative, $\phi_{ABC}=0$ implies $1-I_{5}=0$. According to Proposition 1, since $1-I_{5}$ is a continuous nonnegative function on the pure-state manifold, there exists a pure-state decomposition $\rho=\sum_{j}p_{j}|\psi_{j}\rangle\langle\psi_{j}|$ such that
$\sum_{j}p_{j}\bigl(1-I_{5}(\psi_{j})\bigr)=0\ .$ (32)
Since each term in Eq. (32) is nonnegative, it follows that
$1-I_{5}(\psi_{j})=0$ (33)
for every $j$ with $p_{j}>0$. Hence, every $|{\psi_{j}}\rangle$ is fully product, and thus $\rho$ is fully separable. On the other hand, if $\rho$ is fully separable,
$\tau_{A|B}=\tau_{B|C}=\tau_{C|A}=\tau_{ABC}=1-I_{5}=0\ ,$ (34)
thus $\phi_{ABC}=0$. ∎
Upper bound of $\phi_{ABC}$ for three-qubit pure state.
The upper bound of $\phi_{ABC}$ is $\phi_{ABC}\leq 99/2$, which is saturated by the GHZ-state. See Appendix B for a proof. Hence, the value of $\phi_{ABC}$ also indicates how close a 3-qubit pure state is to the GHZ-state.
2.3 Tripartite Mixed State
The measure $\phi_{ABC}$ is defined for three-qubit states. For general tripartite mixed states, we can construct a criterion for their full separability based on $I_{5}$ and the convex-roof construction [15] as follows.
Proposition 3.
Let [^4] We assign the power $2/3$ in Eq. (35) to make sure that $E_{1-I_{5}}$ is an entanglement measure [42, 45] according to the Theorem 1 in Ref. [17].
$E_{1-I_{5}}(\rho)\equiv\inf_{\{p_{\ell},|{\psi_{\ell}}\rangle\}\in\mathbb{D}(\rho)}\sum_{\ell}p_{\ell}\,\big(1-I_{5}(\psi_{\ell})\big)^{2/3}\ .$ (35)
Then, for any three-qubit mixed state $\rho$,
$E_{1-I_{5}}(\rho)=0\iff\rho\ \text{is fully separable}\ .$ (36)
Proof.
This follows directly from Eq. (19), which gives $I_{5}\leq 1$ and hence $1-I_{5}\geq 0$. ∎
3 Entanglement Measures for Four-Qubit Pure States
Four-qubit pure states provide the simplest but nontrivial setting in which the criteria in Sec. 2 becomes directly useful: tracing out one qubit produces a tripartite mixed state. At the same time, the pure-state system itself already exhibits a rich multipartite structure. A four-qubit pure state modulo local unitary transformations is characterized by eighteen independent real parameters [8], thus a complete description should involve eighteen independent quantities. We identify such a set explicitly in this section. As summarized in Table 1, these eighteen measures consist of six bipartite measures, eight tripartite measures, and four quadripartite measures. We now introduce them and explain how together they form a complete set of entanglement measures for four-qubit pure states. In particular, we introduce how the hyperdeterminant generate these entanglement measures. Finally, we evaluate these entanglement measures for representative states and list the nine entangled ways for 4-qubit pure states.
3.1 Bipartite Entanglement Measures
The two-tangle (2) is an entanglement measure originally defined for two-qubit pure states [3, 24, 49]
$|\Psi\rangle=\sum_{j,k=0}^{1}t_{jk}|jk\rangle\ .$ (37)
The definition (2) can be generalized to finite-dimensional bipartite pure states. If we limit to two-qubit pure states, the two-tangle can also be expressed in alternative forms, such as the square of the two-qubit concurrence [3, 24, 49]
$\tau_{A|B}=C^{2}(\Psi)\ ,\quad\text{with}\ \ C(\Psi)\equiv|\langle\Psi|\sigma_{y}^{\otimes 2}|\Psi^{*}\rangle|\ ,$ (38)
the determinant of the reduced density matrix [3, 24, 49]
$\tau_{A|B}=4\det\rho_{A}\ ,$ (39)
and the hyperdeterminant of the coefficient tensor
$\tau_{A|B}=4|\mathrm{Det}\,\mathrm{T}_{2}|^{2}\ ,$ (40)
where $\mathrm{T}_{2}$ is the $2\times 2$ coefficient tensor of the two-qubit state, with components
$(\mathrm{T}_{2})_{jk}=t_{jk}\ .$ (41)
For a two-qubit mixed state, the convex-roof in Eq. (3) can be computed as in Refs. [24, 49]: Let $\rho_{AB}$ be the density matrix of these two qubits, define the spin flipped density matrix as [24, 49]
$\tilde{\rho}_{AB}\equiv(\sigma_{y}\otimes\sigma_{y})\rho_{AB}^{*}(\sigma_{y}\otimes\sigma_{y})\ .$ (42)
The eigenvalues of $\rho_{AB}\tilde{\rho}_{AB}$ are real and nonnegative. Let $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ be the square roots of these eigenvalues in decreasing order. Then the two-tangle (3) can be obtained using the following formula [24, 49]
$E_{\tau_{A|B}}(\rho)=(\max\{0,\lambda_{1}-\lambda_{2}-\lambda_{3}-\lambda_{4}\})^{2}\ .$ (43)
3.2 Tripartite Entanglement Measures
The three-tangle (6) is a genuine tripartite entanglement measure for three-qubit pure states [12]. It can also be defined as the hyperdeterminant of the coefficient tensor. For a general three-qubit pure state
$|\Psi\rangle=\sum_{j,k,l=0}^{1}t_{jkl}|jkl\rangle\ ,$ (44)
the three-tangle is given by [12]
$\tau_{ABC}=4|\mathrm{Det}\,\mathrm{T}_{3}|=4\left|d_{1}-2d_{2}+4d_{3}\right|\ ,$ (45)
with $\mathrm{T}_{3}$ the $2\times 2\times 2$ coefficient tensor of the three-qubit pure state, $(\mathrm{T}_{3})_{jkl}=t_{jkl}$, and
$\begin{split}d_{1}=&\,t_{000}^{2}t_{111}^{2}+t_{001}^{2}t_{110}^{2}+t_{010}^{2}t_{101}^{2}+t_{100}^{2}t_{011}^{2}\ ,\\ d_{2}=&\,t_{000}t_{111}t_{011}t_{100}+t_{000}t_{111}t_{101}t_{010}+t_{000}t_{111}t_{110}t_{001}\\ &+t_{011}t_{100}t_{101}t_{010}+t_{011}t_{100}t_{110}t_{001}+t_{101}t_{010}t_{110}t_{001}\ ,\\ d_{3}=&\,t_{000}t_{110}t_{101}t_{011}+t_{111}t_{001}t_{010}t_{100}\ .\end{split}$ (46)
Another tripartite quantity is the third-order negativity $I_{5}$ [37]. In Ref. [37], the authors studied the set of local invariants of three-qubit pure states, including two-tangles, the three-tangle (45), and the third-order negativity $I_{5}$ (16). Although $I_{5}$ is not an entanglement measure, combining two- and three-tangles can form an entanglement measure $\phi_{ABC}$ (9[32]. In Sec. 2.2, we have shown that $\phi_{ABC}$ is closely related to $I_{5}$, maintaining the permutational symmetry for the three subsystems in a manifest form (10). We also show that $\phi_{ABC}$ can serve as a necessary and sufficient criterion for the full separability of three-qubit mixed states. Since $\phi_{ABC}$ is an entanglement measure, its role is not limited to providing a yes-or-no criterion for tripartite full separability. The value of $\phi_{ABC}$ itself is also meaningful, so it is natural to study its allowed range and the states that saturate the corresponding bounds, which have been discussed in Sec. 2.2 and Appendix B.
3.3 Quadripartite Entanglement Measures
Given the number of parameters, we still need 4 quadripartite entanglement measures. We start with the four-qubit hyperdeterminant $\Delta$ and show the remaining measures hiding in it. Consider a general four-qubit pure state
$|{\Psi}\rangle=\sum_{j,k,l,m=0}^{1}t_{jklm}|{jklm}\rangle\$(47)
with the following notation
$t_{jklm}:=a_{r}\ ,\quad r:=8j+4k+2l+m\ .$ (48)
We show that the degree-24 hyperdeterminant $\Delta$ [27] can be expressed in terms of objects with four-body permutation symmetry
$\begin{split}\Delta=&\,\mathrm{Det}\,\mathrm{T}_{4}\\ =&\,-\frac{1}{108}\Big(4H^{6}\Pi+6H^{5}\Sigma W-3H^{4}\Sigma^{2}-48H^{3}\Pi W-4H^{3}W^{3}+48H^{2}\Pi\Sigma\\ &\qquad\quad\qquad-60H^{2}\Sigma W^{2}+96H\Sigma^{2}W+64\Pi^{2}+96\Pi W^{2}-32\Sigma^{3}+36W^{4}\Big)\ ,\end{split}$ (49)
where $\Sigma$ and $\Pi$ can be expressed using degree-four invariants $L$, $M$, and $N$:
$\Sigma=L^{2}+M^{2}+N^{2}\ ,\quad\Pi=(L-M)(M-N)(N-L)\ .$ (50)
The degree-four invariants $L$, $M$, and $N$ are defined as
$L=\det\!\begin{pmatrix}a_{0}&a_{4}&a_{8}&a_{12}\\ a_{1}&a_{5}&a_{9}&a_{13}\\ a_{2}&a_{6}&a_{10}&a_{14}\\ a_{3}&a_{7}&a_{11}&a_{15}\end{pmatrix},\ M=\det\!\begin{pmatrix}a_{0}&a_{8}&a_{2}&a_{10}\\ a_{1}&a_{9}&a_{3}&a_{11}\\ a_{4}&a_{12}&a_{6}&a_{14}\\ a_{5}&a_{13}&a_{7}&a_{15}\end{pmatrix},\ N=\det\!\begin{pmatrix}a_{0}&a_{1}&a_{8}&a_{9}\\ a_{2}&a_{3}&a_{10}&a_{11}\\ a_{4}&a_{5}&a_{12}&a_{13}\\ a_{6}&a_{7}&a_{14}&a_{15}\end{pmatrix}$, (51)
and they satisfy $L+M+N=0$, thus only two of them are independent. The quantity $\Sigma$ and $\Pi$ (50) constructed from $L$, $M$, and $N$ are symmetric under permutation among $A$, $B$, $C$, and $D$. Therefore, $\Sigma$ and $\Pi$ are quadripartite measures. It has been shown in Ref. [17] that
$|L|^{2}=\det\rho_{AB}\ ,\quad|M|^{2}=\det\rho_{AC}\ ,\quad|N|^{2}=\det\rho_{AD}\ .$ (52)
For a normalized $4\times 4$ matrix $\rho$, we have the following identity,
$\det\rho=\frac{1}{24}\Big[1-6\operatorname{Tr}(\rho^{2})+3\big(\!\operatorname{Tr}(\rho^{2})\big)^{2}+8\operatorname{Tr}(\rho^{3})-6\operatorname{Tr}(\rho^{4})\Big]\ ,$ (53)
thus we have
$|L|^{2}=\frac{1}{24}\Big[1-6\operatorname{Tr}(\rho_{AB}^{2})+3\big(\!\operatorname{Tr}(\rho_{AB}^{2})\big)^{2}+8\operatorname{Tr}(\rho_{AB}^{3})-6\operatorname{Tr}(\rho_{AB}^{4})\Big]\ ,$ (54)
where $\operatorname{Tr}(\rho_{AB}^{4})$ contains the information about the fourth-order spectral. Therefore, $L$, $M$, and $N$ are independent of the bipartite and tripartite measures that are studied in Secs. 3.1 and 3.2.
The quantity $W$ in Eq. (49) is a degree-6 quantity defined as [38, 27, 30]
$W=D_{xy}+D_{xz}+D_{xw}\ ,$ (55)
with $D_{uv}=\det(B_{uv})$. The $3\times 3$ matrix $B_{uv}$ can be obtained by
$\det\Big(\frac{\partial^{2}A}{\partial u_{i}\partial v_{j}}\Big)=(u_{0}^{2}\quad u_{0}u_{1}\quad u_{1}^{2})B_{uv}\begin{pmatrix}v_{0}^{2}\\ v_{0}v_{1}\\ v_{1}^{2}\end{pmatrix}\$, (56)
with $A$ a quadrilinear form
$A(\mathbf{x},\mathbf{y},\mathbf{z},\mathbf{w})=\sum_{i,j,k,l=0}^{1}t_{ijkl}x_{i}y_{j}z_{k}w_{l}\ .$ (57)
There are some relations between these quantities:
$D_{xy}=D_{zw}\ ,\quad D_{xz}=D_{yw}\ ,\quad D_{xw}=D_{yz}\ ,$ (58)
$HL=D_{xz}-D_{xw}\ ,\quad HM=D_{xw}-D_{xy}\ ,\quad HN=D_{xy}-D_{xz}\ .$ (59)
In Eq. (49), $H$ is a degree-two invariant defined as [27]
$H=a_{0}a_{15}-a_{1}a_{14}-a_{2}a_{13}+a_{3}a_{12}-a_{4}a_{11}+a_{5}a_{10}+a_{6}a_{9}-a_{7}a_{8}\ .$ (60)
We will show that the $H$ in Eq. (60) is precisely the quantity entering the four-tangle (12), by expressing it in terms of the four-fold spin-flip operator $\Omega$ [40, 48]:
$H=\frac{1}{2}\langle\Psi^{*}|\Omega|\Psi\rangle\ ,\quad\Omega:=\sigma_{y}\otimes\sigma_{y}\otimes\sigma_{y}\otimes\sigma_{y}\ .$ (61)
For the mixed state $\rho_{ABCD}$, similar to Eq. (42), define
$\tilde{\rho}_{ABCD}:=\sigma_{y}^{\otimes 4}\rho^{*}_{ABCD}\sigma_{y}^{\otimes 4}\ ,$ (62)
then like Eq. (43), the four-tangle $4|H|^{2}$ can be computed from the squart roots of $\rho_{ABCD}\tilde{\rho}_{ABCD}$ eigenvalues [40, 48]
$4|H|^{2}=C^{2}_{1234}(\rho_{ABCD})=(\max\{0,\lambda_{1}-\lambda_{2}-\cdots-\lambda_{16}\})^{2}\ .$ (63)
To verify Eq. (61), recall that
$\sigma_{y}|{0}\rangle=i|{1}\rangle\ ,\quad\sigma_{y}|{1}\rangle=-i|{0}\rangle\ .$ (64)
Hence, one has
$\Omega|jklm\rangle=i^{4}(-1)^{j+k+l+m}\,|\bar{j}\,\bar{k}\,\bar{l}\,\bar{m}\rangle=(-1)^{j+k+l+m}\,|\bar{j}\,\bar{k}\,\bar{l}\,\bar{m}\rangle\ ,$ (65)
where $\bar{j}:=1-j$. Substituting Eq. (47) into Eq. (61), we obtain
$\langle\Psi^{*}|\Omega|\Psi\rangle$ $=\sum_{j,k,l,m,j^{\prime},k^{\prime},l^{\prime},m^{\prime}}t_{jklm}t_{j^{\prime}k^{\prime}l^{\prime}m^{\prime}}\langle jklm|\Omega|j^{\prime}k^{\prime}l^{\prime}m^{\prime}\rangle$
$=\sum_{j,k,l,m=0}^{1}(-1)^{j+k+l+m}t_{jklm}t_{\bar{j}\,\bar{k}\,\bar{l}\,\bar{m}}\ .$ (66)
Using Eq. (48), Eq. (66) becomes
$\langle\Psi^{*}|\Omega|\Psi\rangle=$ $+a_{0}a_{15}-a_{1}a_{14}-a_{2}a_{13}+a_{3}a_{12}-a_{4}a_{11}+a_{5}a_{10}+a_{6}a_{9}-a_{7}a_{8}$
$+a_{15}a_{0}-a_{14}a_{1}-a_{13}a_{2}+a_{12}a_{3}-a_{11}a_{4}+a_{10}a_{5}+a_{9}a_{6}-a_{8}a_{7}$
$=$  $\,2\big(a_{0}a_{15}-a_{1}a_{14}-a_{2}a_{13}+a_{3}a_{12}-a_{4}a_{11}+a_{5}a_{10}+a_{6}a_{9}-a_{7}a_{8}\big)\ .$ (67)
Combining Eq. (67) with Eq. (61), we arrive at Eq. (60).
Since the four-qubit hyperdeterminant $|\Delta|=|\mathrm{Det}\mathrm{T}_{4}|$ has homogeneous degree $24$, the theorem 1 in Ref. [17] rules it out as an entanglement measure. By contrast, the degree-$4$ quantity $|\Delta|^{1/6}$ is one candidate that satisfies the general monotonicity criterion. Similarly, since $\Sigma$ and $\Pi$ have degrees $8$ and $12$, respectively, the quantities $|\Sigma|^{1/2}$ and $|\Pi|^{1/3}$ are degree-4 entanglement measures. Finally, the four-tangle $4|H|^{2}$ has degree $4$, thus it is also an entanglement measure. [^5] One can also choose $4|H|^{2}$, $|\Sigma|^{1/2}$, $|\Pi|^{1/3}$, and $|W|^{2/3}$ as the four independent quadripartite entanglement measures [38, 27, 30], and here we choose the hyperdeterminant $\Delta$ instead of $W$.
3.4 Examples
We compute all 18 entanglement measures for the following representative states:
  • Generalized GHZ-State
    $|{\widetilde{\mathrm{GHZ}}^{(4)}}\rangle=a|{0000}\rangle+b|{1111}\rangle\ ;$ (68)
  • W-State
    $|{\mathrm{W}}\rangle=\frac{1}{2}\big(|{0001}\rangle+|{0010}\rangle+|{0100}\rangle+|{1000}\rangle\big)\ ;$ (69)
  • Cluster-State
    $|{\mathrm{Cluster}}\rangle=\frac{1}{2}\big(|{0000}\rangle+|{0011}\rangle+|{1100}\rangle-|{1111}\rangle\big)\ ;$ (70)
  • Dicke-State
    $|{\mathrm{Dicke}}\rangle=\frac{1}{\sqrt{6}}\big(|{0011}\rangle+|{0101}\rangle+|{0110}\rangle+|{1001}\rangle+|{1010}\rangle+|{1100}\rangle\big)\ ;$ (71)
  • Generalized Double Bell-State ($|a|^{2}+|b|^{2}=|c|^{2}+|d|^{2}=1$)
    $|{\widetilde{\mathrm{Bell}}}\rangle_{AB}\otimes|{\widetilde{\mathrm{Bell}}}\rangle_{CD}=\big(a|{00}\rangle+b|{11}\rangle\big)\otimes\big(c|{00}\rangle+d|{11}\rangle\big)\ .$ (72)
The results of the entanglement measures for these representative states are:
  • Generalized GHZ-state (68),
    $\begin{split}&\tau_{A|B}=\tau_{A|C}=\tau_{A|D}=\tau_{B|C}=\tau_{B|D}=\tau_{C|D}=0\ ,\\ &\tau_{ABC}=\tau_{ABD}=\tau_{ACD}=\tau_{BCD}=0\ ,\\ &\phi_{ABC}=\phi_{ABD}=\phi_{ABD}=\phi_{BCD}=0\ ,\\ &H=ab\ ,\quad\Sigma=0\ ,\quad\Pi=0\ ,\quad\Delta=0\quad(W=0)\ ;\end{split}$ (73)
  • W-state (69),
    $\begin{split}&\tau_{A|B}=\tau_{A|C}=\tau_{A|D}=\tau_{B|C}=\tau_{B|D}=\tau_{C|D}=1/4\ ,\\ &\tau_{ABC}=\tau_{ABD}=\tau_{ACD}=\tau_{BCD}=0\ ,\\ &\phi_{ABC}=\phi_{ABD}=\phi_{ABD}=\phi_{BCD}=207/8\ ,\\ &H=0\ ,\quad\Sigma=0\ ,\quad\Pi=0\ ,\quad\Delta=0\quad(W=0)\ ;\end{split}$ (74)
  • Cluster-State (70),
    $\begin{split}&\tau_{A|B}=\tau_{A|C}=\tau_{A|D}=\tau_{B|C}=\tau_{B|D}=\tau_{C|D}=0\ ,\\ &\tau_{ABC}=\tau_{ABD}=\tau_{ACD}=\tau_{BCD}=0\ ,\\ &\phi_{ABC}=\phi_{ABD}=\phi_{ABD}=\phi_{BCD}=36\ ,\\ &H=0\ ,\quad\Sigma=\frac{1}{2^{7}}\ ,\quad\Pi=\frac{1}{2^{11}}\ ,\quad\Delta=0\quad(W=0)\ ;\end{split}$ (75)
  • Dicke-State (71):
    $\begin{split}&\tau_{A|B}=\tau_{A|C}=\tau_{A|D}=\tau_{B|C}=\tau_{B|D}=\tau_{C|D}=1/9\ ,\\ &\tau_{ABC}=\tau_{ABD}=\tau_{ACD}=\tau_{BCD}=0\ ,\\ &\phi_{ABC}=\phi_{ABD}=\phi_{ABD}=\phi_{BCD}=173/6\ ,\\ &H=1/2\ ,\quad\Sigma=0\ ,\quad\Pi=0\ ,\quad\Delta=0\quad(W=1/72)\ ;\end{split}$ (76)
  • Generalized Double Bell-State (72):
    $\begin{split}&\tau_{A|B}=4|ab|^{2}\ ,\quad\tau_{C|D}=4|cd|^{2}\\ &\tau_{A|C}=\tau_{A|D}=\tau_{B|C}=\tau_{B|D}=0\ ,\\ &\tau_{ABC}=\tau_{ABD}=\tau_{ACD}=\tau_{BCD}=0\ ,\\ &\phi_{ABC}=\phi_{ABD}=144|ab|^{2}\ ,\quad\phi_{ACD}=\phi_{BCD}=144|cd|^{2}\ ,\\ &H=2abcd\ ,\quad\Sigma=2a^{4}b^{4}c^{4}d^{4}\ ,\quad\Pi=-2a^{6}b^{6}c^{6}d^{6}\ ,\quad\Delta=0\quad(W=2a^{3}b^{3}c^{3}d^{3})\ .\end{split}$ (77)
From the generalized GHZ-state, we find that only the four-tangle survives and can quantify quantum entanglement. This observation will be discussed further in Sec. 4. The hyperdeterminant $\Delta$ always vanishes for these states. Therefore, we are interested in exploring further the all entangled classes of four-qubit pure states [^6] Note that the states listed here have not been normalized. [43]:
$\begin{split}|{G_{abcd}^{1}}\rangle=&\,\frac{a+d}{2}\big(|{0000}\rangle+|{1111}\rangle\big)+\frac{a-d}{2}\big(|{0011}\rangle+|{1100}\rangle\big)\\ &+\frac{b+c}{2}\big(|{0101}\rangle+|{1010}\rangle\big)+\frac{b-c}{2}\big(|{0110}\rangle+|{1001}\rangle\big)\ ;\end{split}$ (78)
$\begin{split}|{G_{abc}^{2}}\rangle=&\,\frac{a+b}{2}\big(|{0000}\rangle+|{1111}\rangle\big)+\frac{a-b}{2}\big(|{0011}\rangle+|{1100}\rangle\big)\\ &+c\big(|{0101}\rangle+|{1010}\rangle\big)+|{0110}\rangle\ ;\end{split}$ (79)
$|{G_{ab}^{3}}\rangle=a\big(|{0000}\rangle+|{1111}\rangle\big)+b\big(|{0101}\rangle+|{1010}\rangle\big)+|{0110}\rangle+|{0011}\rangle\ ;$ (80)
$\begin{split}|{G_{ab}^{4}}\rangle=&\,a\big(|{0000}\rangle+|{1111}\rangle\big)+\frac{a+b}{2}\big(|{0101}\rangle+|{1010}\rangle\big)+\frac{a-b}{2}\big(|{0110}\rangle+|{1001}\rangle\big)\\ &+\frac{i}{\sqrt{2}}\big(|{0001}\rangle+|{0010}\rangle+|{0111}\rangle+|{1011}\rangle\big)\ ;\end{split}$ (81)
$|{G_{a}^{5}}\rangle=a\big(|{0000}\rangle+|{0101}\rangle+|{1010}\rangle+|{1111}\rangle\big)+i|{0001}\rangle+|{0110}\rangle-i|{1011}\rangle\ ;$ (82)
$|{G_{a}^{6}}\rangle=a\big(|{0000}\rangle+|{1111}\rangle\big)+|{0011}\rangle+|{0101}\rangle+|{0110}\rangle\ ;$ (83)
$|{G^{7}}\rangle=|{0000}\rangle+|{0101}\rangle+|{1000}\rangle+|{1110}\rangle\ ;$ (84)
$|{G^{8}}\rangle=|{0000}\rangle+|{1011}\rangle+|{1101}\rangle+|{1110}\rangle\ ;$ (85)
$|{G^{9}}\rangle=|{0000}\rangle+|{0111}\rangle\ .$ (86)
The results of the quadripartite entanglement measures are listed in Table 2, and $\Delta$ is only non-zero for $G_{abcd}^{1}$ [27]. Hence, $\Delta$ characterizes a very specific type of quadripartite entanglement.
State $H$ $L$ $M$ $D_{xw}$ $\Delta$ $G_{abcd}^{1}$ (78) $\frac{a^{2}+b^{2}+c^{2}+d^{2}}{2}$ abcdabcd $(\frac{c-d}{2})^{2}\!-\!(\frac{a-b}{2})^{2}][(\frac{a+b}{2})^{2}\!-\!(\frac{c+d}{2})^{2}]$ $\frac{(ad-bc)(bd-ac)(ab-cd)}{4}$ $\frac{V(a^{2},b^{2},c^{2},d^{2})}{256}$ $G_{abc}^{2}$ (79) $\frac{a^{2}+b^{2}+2c^{2}}{2}$ $abc^{2}$ $c^{2}-(\frac{a+b}{2})^{2}](\frac{a-b}{2})^{2}$ $\frac{c^{2}(a-b)^{2}(c^{2}-ab)}{4}$ 0 $G_{ab}^{3}$ (80) $a^{2}+b^{2}$ $a^{2}b^{2}$ 0 0 0 $G_{ab}^{4}$ (81) $\frac{3a^{2}+b^{2}}{2}$ $a^{3}b$ $a^{2}-(\frac{a+b}{2})^{2}](\frac{a-b}{2})^{2}$ $\frac{a^{3}(a-b)^{3}}{4}$ 0 $G_{a}^{5}$ (82) $2a^{2}$ $a^{4}$ 0 0 0 $G_{a}^{6}$ (83) $a^{2}$ 0 0 0 0 $G^{7}$ (84) 0 0 0 0 0 $G^{8}$ (85) 0 0 0 0 0 $G^{9}$ (86) 0 0 0 0 0
Table 2: Quadripartite entanglement measures for all different entangled classes [27]. The Vandermonde determinant is $V(a^{2},b^{2},c^{2},d^{2})=(a^{2}-b^{2})^{2}(a^{2}-c^{2})^{2}(a^{2}-d^{2})^{2}(b^{2}-c^{2})^{2}(b^{2}-d^{2})^{2}(c^{2}-d^{2})^{2}$.
4 Independent Entanglement Measure
The next question is whether one of these measures can exist independently, namely, whether we can keep a single entanglement measure nonvanishing while turning off all the others. Among the eighteen measures, the four quantities $\phi_{ABC}$, $\phi_{ABD}$, $\phi_{ACD}$, and $\phi_{BCD}$ play a distinguished role. As shown in Sec. 2.2, each of them provides a necessary and sufficient criterion for the full separability of the corresponding three-qubit mixed state. Therefore, if one wishes to turn off all four $\phi$’s, one is forced to require all four three-qubit reduced density matrices to be fully separable. The Proposition 1 of Ref. [39] shows that: For a $\mathtt{q}$-qubit pure state, suppose that, for every qubit $X$, the $(\mathtt{q}-1)$-qubit reduced density matrix $\rho_{\bar{X}}=\operatorname{Tr}_{X}|\Psi\rangle\langle\Psi|$ is fully separable, then $|\Psi\rangle$ is LU-equivalent to an $\mathtt{q}$-qubit generalized GHZ-state
$|{\widetilde{\mathrm{GHZ}}^{(\mathtt{q})}}\rangle=a|0\rangle^{\otimes\mathtt{q}}+b|1\rangle^{\otimes\mathtt{q}}\ ,$ (87)
and conversely, any state of the form (87) satisfies the above property. [^7] We will extend this proposition to general $\mathtt{q}$-qudit pure states in Sec. 5.2.
For $\mathtt{q}=4$, this implies that any four-qubit pure state with
$\phi_{ABC}=\phi_{ABD}=\phi_{ACD}=\phi_{BCD}=0$ (88)
is LU equivalent to a generalized GHZ-type four-qubit state
$|{\widetilde{\mathrm{GHZ}}^{(4)}}\rangle=a|{0000}\rangle+b|{1111}\rangle\ ,$ (89)
whose 18 entanglement measures are given in Eq. (73). One can see that, except for $H=ab$, all the other 17 measures vanish. This observation already shows that among the eighteen entanglement measures, four-tangle can survive independently, and this occurs uniquely for the generalized GHZ-state. We now show that this is in fact the only possibility, by showing that $\phi_{ABC}$ cannot exist independently:
Proposition 4.
There is no four-qubit pure state $|\Psi\rangle\in(\mathbb{C}^{2})^{\otimes 4}$ such that
  1. (i)
    $\rho_{BCD}=\operatorname{Tr}_{A}|\Psi\rangle\langle\Psi|$ is fully separable,
  2. (ii)
    $\rho_{ACD}=\operatorname{Tr}_{B}|\Psi\rangle\langle\Psi|$ is fully separable,
  3. (iii)
    $\rho_{ABD}=\operatorname{Tr}_{C}|\Psi\rangle\langle\Psi|$ is fully separable,
  4. (iv)
    $\rho_{ABC}=\operatorname{Tr}_{D}|\Psi\rangle\langle\Psi|$ is not fully separable,
  5. (v)
    $\tau_{ABC}=\tau_{ABD}=\tau_{ACD}=\tau_{BCD}=0$.
Proof.
Assume, for contradiction, that such a state $|\Psi\rangle$ exists. By assumption (i), $\rho_{BCD}$ is a fully separable three-qubit state with rank at most 2. Since $\rho_{ABC}$ is not fully separable, we further require that $\operatorname{rank}(\rho_{BCD})=2$, otherwise $|{\Psi}\rangle$ is a fully product state. Therefore, one can write
$|\Psi\rangle=|a_{0}\rangle|b_{0}\rangle|c_{0}\rangle|d_{0}\rangle+|a_{1}\rangle|b_{1}\rangle|c_{1}\rangle|d_{1}\rangle\ .$ (90)
We now prove a simple lemma.
Lemma 1.
Let $|x_{0}\rangle,|x_{1}\rangle$ be linearly independent vectors in $\mathcal{H}_{X}$, and $|y_{0}\rangle,|y_{1}\rangle$ be linearly independent vectors in $\mathcal{H}_{Y}$. Consider a positive semidefinite operator $\sigma$
$\sigma=\mu_{0}|x_{0}y_{0}\rangle\langle x_{0}y_{0}|+\mu_{1}|x_{1}y_{1}\rangle\langle x_{1}y_{1}|+s\,|x_{0}y_{0}\rangle\langle x_{1}y_{1}|+s^{*}\,|x_{1}y_{1}\rangle\langle x_{0}y_{0}|\ ,$ (91)
with $\mu_{0},\mu_{1}\geq 0$. Then $\sigma$ is separable across $X|Y$ if and only if $s=0$.
Proof.
By invertible local transformations, $\sigma$ is equivalent to
$\mu_{0}|00\rangle\langle 00|+\mu_{1}|11\rangle\langle 11|+s^{\prime}|00\rangle\langle 11|+s^{\prime*}|11\rangle\langle 00|\ ,$ (92)
where $s^{\prime}=0$ if and only if $s=0$. Its partial transpose has eigenvalues $\{\mu_{0},\mu_{1},|s^{\prime}|,-|s^{\prime}|\}$. Hence, it is PPT if and only if $s^{\prime}=0$. By the $2\times 2$ PPT criterion [35, 26], this is equivalent to separability. ∎
Tracing out subsystem $B$ in Eq. (90) and applying Lemma 1 with assumption (i), one can obtain $\langle b_{1}|b_{0}\rangle=0$. Similarly, tracing out subsystem $C$ or subsystem $A$ gives $\langle c_{1}|c_{0}\rangle=0$, $\langle a_{1}|a_{0}\rangle=0$, respectively. Therefore, up to local unitaries,
$|\Psi\rangle=|000\rangle_{ABC}|d_{0}\rangle+|111\rangle_{ABC}|d_{1}\rangle\ .$ (93)
Tracing out subsystem $D$ in Eq. (93), we find
$\rho_{ABC}=\langle d_{0}|d_{0}\rangle|000\rangle\langle 000|+\langle d_{1}|d_{1}\rangle|111\rangle\langle 111|+\langle d_{1}|d_{0}\rangle|000\rangle\langle 111|+\langle d_{0}|d_{1}\rangle|111\rangle\langle 000|\ ,$ (94)
If $\langle d_{0}|d_{1}\rangle=0$, then $\rho_{ABC}$ in Eq. (94) is fully separable, contradicting assumption (iv). Hence, we must have $\langle d_{0}|d_{1}\rangle\neq 0$, then $\rho_{ABC}$ is a rank-$2$ GHZ-type state supported on
$\mathrm{span}\{|000\rangle,|111\rangle\}\ ,$ (95)
with nonzero coherence between $|000\rangle$ and $|111\rangle$. In this sense, $\phi_{ABC}$ is not zero. Such a state also has a nonvanishing three-tangle. This contradicts assumption (v). Therefore, assumptions (i)–(v) cannot be satisfied simultaneously. ∎
Our result shows that the four-tangle $4|H|^{2}$ [40, 48] is the only one among the eighteen entanglement measures that can exist independently. To quantify the quantum entanglement, a state should only involve one entanglement measure. Otherwise, the quantification remains ambiguous under the current understanding. Hence, our results show that the generalized GHZ-state is the only state for which we can quantify quantum entanglement without ambiguity using the four-tangle.
5 Generalization to Multipartite Systems
The discussion so far has focused on tripartite full separability and the special role in its application to the four-qubit pure states. However, the underlying logic is not restricted to these low-party and few-qubit examples. The criterion based on third-order negativity could be extended to a more general system (arbitrary finite-dimensional multipartite systems). It is therefore natural to ask to what extent the separability criteria encountered above admit a general multipartite formulation. In this section, we show that the answer is affirmative.
We first generalize the tripartite construction and obtain necessary and sufficient criteria for full separability in general finite-dimensional multipartite systems. We then show that the same framework also leads to a multipartite rigidity statement: for a $\mathtt{q}$-partite pure state, if all its $(\mathtt{q}-1)$-partite reduced density matrices are fully separable, the state must be LU-equivalent to a generalized GHZ-state. This result is particularly interesting for extending to field theory. Because the local QFT does not have a separable state for the spatial partition [47], we cannot have a generalized GHZ-state for it. In other words, the qudit state’s quantum entanglement should correspond to the mode-wise entanglement on the field-theory side.
5.1 Replica Criterion for Fully Separable Qudit States
Proposition 1 shows that the condition
$I_{5}=1$ (96)
provides a criterion for tripartite fully product states. Since $I_{5}$ admits a replica construction (16), it is natural to ask whether, for multipartite pure states, one can construct analogous quantities from the replica that can serve as criteria for multipartite fully product states.
We consider $\mathcal{Z}_{\Omega_{1,\cdots,\mathtt{q}}}^{(\mathtt{q})}$ (100), called multi-invariant [19, 20], to generalize Proposition 1 to $\mathtt{q}$-partite setting: $|\mathcal{Z}_{\Omega_{1,\cdots,\mathtt{q}}}^{(\mathtt{q})}|=1$ can serve as a criterion for $\mathtt{q}$-partite fully product states (Proposition 5). Let
$|\Psi\rangle=\sum_{i_{1}=0}^{d_{1}-1}\cdots\sum_{i_{\mathtt{q}}=0}^{d_{\mathtt{q}}-1}t_{i_{1}\cdots i_{\mathtt{q}}}|i_{1}\cdots i_{\mathtt{q}}\rangle\in\mathbb{C}^{d_{1}}\otimes\cdots\otimes\mathbb{C}^{d_{\mathtt{q}}}$ (97)
be a normalized $\mathtt{q}$-qudit pure state. Let $N$ be the number of replicas, and let
$\Omega_{r}\in S_{N}\ \ \text{for}\ \ r=1,\dots,\mathtt{q}\ ,$ (98)
with
$\Omega_{r}\neq\Omega_{s}\ \ \text{for all}\ \ 1\leq r<s\leq\mathtt{q}\ .$ (99)
Define
$\mathcal{Z}_{\Omega_{1,\cdots,\mathtt{q}}}^{(\mathtt{q})}(\Psi)\equiv\bigl\langle\Psi^{\otimes N}\big|\Omega_{1}\Omega_{2}\cdots\Omega_{\mathtt{q}}\big|\Psi^{\otimes N}\bigr\rangle\ ,$ (100)
where $\Omega_{r}$ acts on the $r$-th subsystem across the $N$ replicas. Note that the quantity in Eq. (100) is invariant under $\Omega_{1,\cdots,\mathtt{q}}\to\Omega^{\prime}\Omega_{1,\cdots,\mathtt{q}}$, thus one can always set $\Omega_{1}=\mathrm{id}$ by taking $\Omega^{\prime}=(\Omega_{1})^{-1}$.
Proposition 5.
For a normalized pure state $|\Psi\rangle\in\mathbb{C}^{d_{1}}\otimes\cdots\otimes\mathbb{C}^{d_{\mathtt{q}}}$,
$\big|\mathcal{Z}_{\Omega_{1,\cdots,\mathtt{q}}}^{(\mathtt{q})}(\Psi)\big|=1\quad\Longleftrightarrow\quad|\Psi\rangle\ \text{is fully product}\ .$ (101)
Proof.
The direction “fully product $\Rightarrow\big|\mathcal{Z}_{\Omega_{1,\cdots,\mathtt{q}}}^{(\mathtt{q})}(\Psi)\big|=1$” is immediate. We now prove the converse. Let
$\Omega:=\Omega_{1}\Omega_{2}\cdots\Omega_{\mathtt{q}}\ .$ (102)
Since $\Omega$ is unitary, one immediately has
$\big|\mathcal{Z}_{\Omega_{1,\cdots,\mathtt{q}}}^{(\mathtt{q})}(\Psi)\big|\leq 1\ .$ (103)
If $|\mathcal{Z}_{\Omega_{1,\cdots,\mathtt{q}}}^{(\mathtt{q})}(\Psi)|=1$, then $|\Psi\rangle^{\otimes N}$ is an eigenvector of $\Omega$ with eigenvalue $e^{i\theta}$, namely
$\Omega|\Psi\rangle^{\otimes N}=e^{i\theta}|\Psi\rangle^{\otimes N}\ ,\quad\theta\in[0,2\pi)\ .$ (104)
We first rewrite Eq. (104) as an identity for the coefficients $t_{i_{1}\cdots i_{\mathtt{q}}}$. For each replica $x\in\{1,\dots,N\}$, we choose local indices
$i_{r}(x)\in\{0,\dots,d_{r}-1\}\ ,\quad r=1,\dots,\mathtt{q}\ .$ (105)
Comparing the coefficient of the corresponding computational basis vector on both sides of Eq. (104), we obtain
$\prod_{x=1}^{N}t_{\,i_{1}(\Omega_{1}(x)),\,i_{2}(\Omega_{2}(x)),\,\dots,\,i_{\mathtt{q}}(\Omega_{\mathtt{q}}(x))}=e^{i\theta}\prod_{x=1}^{N}t_{\,i_{1}(x),\,i_{2}(x),\,\cdots,\,i_{\mathtt{q}}(x)}\ .$ (106)
We next show that $\theta$ must be zero. Since $|\Psi\rangle$ is normalized, there exists at least one choice of indices
$(i_{1},\dots,i_{\mathtt{q}})=(\mu_{1},\dots,\mu_{\mathtt{q}})$ (107)
such that $t_{\mu_{1}\cdots\mu_{\mathtt{q}}}\neq 0$. Now we apply the assignment (107) for all $x=1,\dots,N$, then Eq. (106) gives
$\bigl(t_{\mu_{1}\cdots\mu_{\mathtt{q}}}\bigr)^{N}=e^{i\theta}\bigl(t_{\mu_{1}\cdots\mu_{\mathtt{q}}}\bigr)^{N}\ .$ (108)
Since $t_{\mu_{1}\cdots\mu_{\mathtt{q}}}\neq 0$, this implies $e^{i\theta}=1$ and $\theta=0$.
We finally show that every $2\times 2$ submatrix of every two-index slice of $T$ vanishes, with $(T)_{i_{1}i_{2}\cdots i_{\mathtt{q}}}=t_{i_{1}i_{2}\cdots i_{\mathtt{q}}}$ the $d_{1}\times d_{2}\times\cdots\times d_{\mathtt{q}}$ coefficient tensor of the $\mathtt{q}$-qudit pure state (97). Consider a slice involving the $r$-th and $s$-th indices, where $1\leq r<s\leq\mathtt{q}$. Since $\Omega_{r}\neq\Omega_{s}$, choose $x_{*}$ such that
$\Omega_{r}^{-1}(x_{*})\neq\Omega_{s}^{-1}(x_{*})\ .$ (109)
Fix all indices except $(i_{r},i_{s})$, and choose two values
$\alpha_{r}\neq\beta_{r}\in\{0,\dots,d_{r}-1\}\ ,\quad\alpha_{s}\neq\beta_{s}\in\{0,\dots,d_{s}-1\}\ .$ (110)
Let
$\begin{split}&A:=t_{\mu_{1}\dots\mu_{r-1}\alpha_{r}\mu_{r+1}\cdots\mu_{s-1}\alpha_{s}\mu_{s+1}\cdots\mu_{\mathtt{q}}}\ ,\quad B:=t_{\mu_{1}\dots\mu_{r-1}\alpha_{r}\mu_{r+1}\cdots\mu_{s-1}\beta_{s}\mu_{s+1}\cdots\mu_{\mathtt{q}}}\ ,\\ &C:=t_{\mu_{1}\dots\mu_{r-1}\beta_{r}\mu_{r+1}\cdots\mu_{s-1}\alpha_{s}\mu_{s+1}\cdots\mu_{\mathtt{q}}}\ ,\quad D:=t_{\mu_{1}\dots\mu_{r-1}\beta_{r}\mu_{r+1}\cdots\mu_{s-1}\beta_{s}\mu_{s+1}\cdots\mu_{\mathtt{q}}}\ ,\end{split}$ (111)
where $\mu_{\dots}$ denotes the fixed values of the remaining $\mathtt{q}-2$ indices. We claim that
BC=AD.BC=AD\ . (112)
If all four entries vanish, there is nothing to prove. Otherwise, after exchanging $\alpha_{r}\leftrightarrow\beta_{r}$ and/or $\alpha_{s}\leftrightarrow\beta_{s}$ if necessary, we can assume $D\neq 0$. Choose a copy labeled by $x_{*}\in\{1,\dots,N\}$. Apply the following assignment
$\text{at }x_{*}:\ (i_{1,\dots,r-1},i_{r},i_{r+1,\dots,s-1},i_{s},i_{s+1,\dots,\mathtt{q}})=(\mu_{1,\dots,r-1},\alpha_{r},\mu_{r+1,\cdots,s-1},\alpha_{s},\mu_{s+1,\cdots,\mathtt{q}})\ ,$
$\text{at }x\neq x_{*}:\ (i_{1,\dots,r-1},i_{r},i_{r+1,\dots,s-1},i_{s},i_{s+1,\dots,\mathtt{q}})=(\mu_{1,\dots,r-1},\beta_{r},\mu_{r+1,\cdots,s-1},\beta_{s},\mu_{s+1,\cdots,\mathtt{q}})$
to Eq. (106), then the LHS of Eq. (106) is $BCD^{N-2}$, while the RHS is $AD^{N-1}$. [^8] Recall that we have already proved that $\theta=0$ in Eq. (106). If one sets $\alpha_{r}=\beta_{r}$, this gives the identity $AB=BA$ instead of Eq. (112) as we require. Since $D\neq 0$, we obtain Eq. (112). Therefore, the determinant of every $2\times 2$ submatrix of every two-index slice of $T$ vanishes. By the standard characterization of rank-one tensors in terms of flattenings [^9] A tensor $T$ is rank one if and only if every flattening of $T$ has rank at most one, equivalently, all $2\times 2$ submatrices of all flattenings of $T$ vanish. , this implies that $T$ is rank one, i.e.,
$t_{i_{1}\cdots i_{\mathtt{q}}}=u^{(1)}_{i_{1}}\cdots u^{(\mathtt{q})}_{i_{\mathtt{q}}}$ (113)
for some vectors $u^{(r)}\in\mathbb{C}^{d_{r}}$. Therefore, we get
$|\Psi\rangle=\Biggl(\sum_{i_{1}=0}^{d_{1}-1}u^{(1)}_{i_{1}}|i_{1}\rangle\Biggr)\otimes\cdots\otimes\Biggl(\sum_{i_{\mathtt{q}}=0}^{d_{\mathtt{q}}-1}u^{(\mathtt{q})}_{i_{\mathtt{q}}}|i_{\mathtt{q}}\rangle\Biggr)\ ,$ (114)
so $|\Psi\rangle$ is fully product. ∎
The key point of this construction is that the extremal condition $|\mathcal{Z}|=1$ turns a replica invariant into a rigid fixed-vector constraint on the replicated state, i.e., Eq. (104) with $\theta=0$. This immediately yields polynomial relations among the coefficients of $|\Psi\rangle$, from which one shows that every two-index slice has vanishing determinant. Hence, the coefficient tensor is forced to be rank one, so $|\Psi\rangle$ is fully product. Since the symmetric group $S_{N}$ has $N!$ elements, in principle one needs at least the smallest integer $N$ such that $N!\geq\mathtt{q}$ replicas to determine whether a $\mathtt{q}$-partite state is fully product.
Several important examples are listed below:
  • Bipartite $n$-th Rényi entropy: $N=n$, $\Omega_{1}=\mathrm{id}$, and $\Omega_{2}=(12\cdots n)$. The corresponding replica partite function $\mathcal{Z}_{n}^{(2)}$ is real number.
  • Third-order negativity (16): $N=3$, $\Omega_{1}=\mathrm{id}$, $\Omega_{2}=(123)$, and $\Omega_{3}=(132)$. The corresponding replica partite function $\mathcal{Z}_{\Omega_{1,2,3}}^{(3)}$ is real number.
  • $\mathtt{q}$-partite $2$-nd Rényi multi-entropy: there are $N=2^{\mathtt{q}-1}$ replicas, and we label these replicas by vectors
    $x=(x_{2},\dots,x_{\mathtt{q}})\in\mathbb{Z}_{2}^{\mathtt{q}-1}\ .$ (115)
    For each $r=2,\dots,\mathtt{q}$, let $\Omega_{r}$ be the permutation operator that flips the $r$-th hypercube coordinate,
    $\Omega_{r}:\ x\mapsto x+e_{r}\ ,$ (116)
    where $e_{r}$ is the unit vector in the $r$-th direction. The corresponding replica partite function
    $\mathcal{Z}_{2}^{(\mathtt{q})}(\Psi)\equiv\bigl\langle\Psi^{\otimes N}\big|\Omega_{2}\Omega_{3}\cdots\Omega_{\mathtt{q}}\big|\Psi^{\otimes N}\bigr\rangle$ (117)
    is a real number. See Figure 1 for the graphical notation of the wavefunction, reduced density matrix, and $\mathcal{Z}_{2}^{(3)}$ from the contractions of reduced density matrices.
The three examples above share two features that make them especially natural: their corresponding replica partition functions are directly real-valued, and they are invariant under arbitrary permutations of the parties. Similar to Sec. 2.3, Proposition 5 can also be generalized to mixed states using the convex-roof construction.
Proposition 6.
For every normalized $\mathtt{q}$-qudit mixed state, let [^10] According to the Theorem 1 of Ref. [17], we assign the power $2/N$ in Eq. (118) to make sure that $E_{1-|\mathcal{Z}_{\Omega_{1,\cdots,\mathtt{q}}}^{(\mathtt{q})}|}$ is an entanglement measure.
$E_{1-|\mathcal{Z}_{\Omega_{1,\cdots,\mathtt{q}}}^{(\mathtt{q})}|}(\rho)\equiv\inf_{\{p_{\ell},|{\psi_{\ell}}\rangle\}\in\mathbb{D(\rho)}}\sum_{\ell}p_{\ell}\Big(1-\big|\mathcal{Z}_{\Omega_{1,\cdots,\mathtt{q}}}^{(\mathtt{q})}(\psi_{\ell})\big|\Big)^{2/N}\ ,$ (118)
then,
$E_{1-|\mathcal{Z}_{\Omega_{1,\cdots,\mathtt{q}}}^{(\mathtt{q})}|}(\rho)=0\iff\rho\ \text{is fully separable}\ .$ (119)
Proof.
This follows directly from Eq. (103) and Proposition 5. ∎
Figure 1: Graphical notation for the tripartite wavefunction $\Psi$, its conjugate $\Psi^{*}$, the reduced density matrix $\rho_{BC}$, and the construction of $\mathcal{Z}^{(\mathtt{q})}_{2}$ in the special case $\mathtt{q}=3$.
5.2 Subsystem Separability and Generalized GHZ-States
Here we generalized Proposition 1 of Ref. [39] to finite-dimensional $\mathtt{q}$-partite states. Namely, for a $\mathtt{q}$-partite pure state, requiring all $(\mathtt{q}-1)$-partite reduced density matrices to be fully separable turns out to be an extremely restrictive condition, which forces the state to be LU-equivalent to a generalized GHZ-state. We formulate this precisely in the following proposition.
Proposition 7.
Let
$|\Psi\rangle\in\mathcal{H}_{1}\otimes\mathcal{H}_{2}\otimes\cdots\otimes\mathcal{H}_{\mathtt{q}}$ (120)
be a $\mathtt{q}$-qudit pure state, where all Hilbert spaces are finite-dimensional. Suppose that for every qudit $X$, the $(\mathtt{q}-1)$-qudit reduced density matrix
$\rho_{\bar{X}}=\operatorname{Tr}_{X}|\Psi\rangle\langle\Psi|$ (121)
is fully separable. Then $|\Psi\rangle$ is LU-equivalent to a rank-$R$ $\mathtt{q}$-qudit generalized GHZ-state
$|{\widetilde{\mathrm{GHZ}}_{R}^{(\mathtt{q})}}\rangle=\sum_{j=0}^{R-1}\lambda_{j}|j\rangle^{\otimes\mathtt{q}}\ ,$ (122)
where $\sum_{j}|\lambda_{j}|^{2}=1$ and
$R\leq\min\{\dim\mathcal{H}_{1},\dim\mathcal{H}_{2},\dots,\dim\mathcal{H}_{\mathtt{q}}\}\ .$ (123)
Conversely, any state of the form (122) satisfies the above property.
Proof.
Here we present the proof only for the four-partite case; the extension to general $\mathtt{q}$ is straightforward. The converse is immediate. We now prove the nontrivial direction.
Let us relabel the four parties by $A$, $B$, $C$, and $D$. Since $\rho_{ABC}$ is fully separable, its support is spanned by fully product vectors. Let
$R=\operatorname{rank}(\rho_{ABC})=\operatorname{rank}(\rho_{D})\ .$ (124)
Choose $R$ linearly independent fully product vectors
$|a_{j}b_{j}c_{j}\rangle:=|a_{j}\rangle\otimes|b_{j}\rangle\otimes|c_{j}\rangle\ ,\quad j=1,\dots,R$ (125)
spanning $\operatorname{supp}(\rho_{ABC})$. Since $|\Psi\rangle$ is a purification of $\rho_{ABC}$ and $\operatorname{rank}(\rho_{D})=R$, we can write it as
$|\Psi\rangle=\sum_{j=1}^{R}|a_{j}b_{j}c_{j}\rangle|d_{j}\rangle\ ,$ (126)
where the vectors
$\{|d_{j}\rangle\}_{j=1}^{R}\subset\mathcal{H}_{D}$ (127)
are linearly independent.
Tracing out $A$ in Eq. (126), we obtain
$\rho_{BCD}=\sum_{j,k=1}^{R}\langle a_{k}|a_{j}\rangle\,|b_{j}c_{j}d_{j}\rangle\langle b_{k}c_{k}d_{k}|\ .$ (128)
Since $\rho_{BCD}$ is fully separable, it is separable across $BC|D$. Since the vectors $\{|d_{k}\rangle\}_{j=1}^{R}$ are linearly independent, there exists a linear map
$M_{D}:\ \mathcal{H}_{D}\to\mathbb{C}^{2}$ (129)
such that
$M_{D}|d_{j}\rangle=|0\rangle\ ,\quad M_{D}|d_{k}\rangle=|1\rangle\ ,\quad M_{D}|d_{l}\rangle=0$ (130)
for $l\neq j,k$. Applying the local filter $I_{BC}\otimes M_{D}$ to $\rho_{BCD}$, we obtain
$\tilde{\rho}_{BCD}=(I_{BC}\otimes M_{D})\rho_{BCD}(I_{BC}\otimes M_{D}^{\dagger})\ ,$ (131)
which is still separable across $BC|D$. Using Eq. (128), we find
$\tilde{\rho}_{BCD}$ $=\langle a_{j}|a_{j}\rangle\,|b_{j}c_{j}\rangle\langle b_{j}c_{j}|\otimes|0\rangle\langle 0|+\langle a_{k}|a_{k}\rangle\,|b_{k}c_{k}\rangle\langle b_{k}c_{k}|\otimes|1\rangle\langle 1|$
$\quad+\langle a_{k}|a_{j}\rangle\,|b_{j}c_{j}\rangle\langle b_{k}c_{k}|\otimes|0\rangle\langle 1|+\langle a_{j}|a_{k}\rangle\,|b_{k}c_{k}\rangle\langle b_{j}c_{j}|\otimes|1\rangle\langle 0|\ .$ (132)
If $|b_{j}c_{j}\rangle$ and $|b_{k}c_{k}\rangle$ are linearly independent, then $\tilde{\rho}_{BCD}$ is of the form (91) across $BC|D$. Hence, by Lemma 1,
$\langle a_{j}|a_{k}\rangle\neq 0\quad\Longrightarrow\quad|b_{k}c_{k}\rangle\propto|b_{j}c_{j}\rangle\quad\Longrightarrow\quad|b_{k}\rangle\propto|b_{j}\rangle\ ,\ |c_{k}\rangle\propto|c_{j}\rangle\ .$ (133)
Cyclically permuting $A,B,C$, we likewise obtain
$\langle b_{j}|b_{k}\rangle\neq 0\quad\Longrightarrow\quad|c_{k}\rangle\propto|c_{j}\rangle\ ,\ |a_{k}\rangle\propto|a_{j}\rangle\ ,$ (134)
and
$\langle c_{j}|c_{k}\rangle\neq 0\quad\Longrightarrow\quad|a_{k}\rangle\propto|a_{j}\rangle\ ,\ |b_{k}\rangle\propto|b_{j}\rangle\ .$ (135)
We now show that $\{|a_{j}\rangle\}$, $\{|b_{j}\rangle\}$, and $\{|c_{j}\rangle\}$ are pairwise orthogonal. If $\langle a_{j}|a_{k}\rangle\neq 0$, then Eq. (133) implies $|b_{k}\rangle\propto|b_{j}\rangle$ and $|c_{k}\rangle\propto|c_{j}\rangle$, thus $\langle b_{j}|b_{k}\rangle\neq 0$, and then Eq. (134) gives $|a_{k}\rangle\propto|a_{j}\rangle$. Thus,
$|a_{k}b_{k}c_{k}\rangle\propto|a_{j}b_{j}c_{j}\rangle\ ,$ (136)
contradicting the linear independence of $\{|a_{j}b_{j}c_{j}\rangle\}_{j=1}^{R}$. By the same logic, starting from $\langle b_{j}|b_{k}\rangle\neq 0$ or $\langle c_{j}|c_{k}\rangle\neq 0$, we also get $\langle b_{j}|b_{k}\rangle=\langle c_{j}|c_{k}\rangle=0$ for $j\neq k$. Therefore, we get
$\langle a_{j}|a_{k}\rangle=\langle b_{j}|b_{k}\rangle=\langle c_{j}|c_{k}\rangle=0\qquad(j\neq k)\ .$ (137)
Tracing out $D$ in Eq. (126), we get
$\rho_{ABC}=\sum_{j,k=1}^{R}\langle d_{k}|d_{j}\rangle\,|a_{j}b_{j}c_{j}\rangle\langle a_{k}b_{k}c_{k}|\ .$ (138)
Since $\rho_{ABC}$ is fully separable, it is separable across $A|BC$. For any $j\neq k$, projecting $\rho_{ABC}$ onto
$\operatorname{Span}\{|a_{j}\rangle,|a_{k}\rangle\}\otimes\operatorname{Span}\{|b_{j}c_{j}\rangle,|b_{k}c_{k}\rangle\}$ (139)
and using Lemma 1, we find
$\langle d_{k}|d_{j}\rangle=0\qquad(k\neq j)\ .$ (140)
In summary, each of the four families
$\{|a_{j}\rangle\}_{j=1}^{R}\ ,\quad\{|b_{j}\rangle\}_{j=1}^{R}\ ,\quad\{|c_{j}\rangle\}_{j=1}^{R}\ ,\quad\{|d_{j}\rangle\}_{j=1}^{R}$ (141)
is pairwise orthogonal. Therefore, there exist local unitaries sending them to the standard basis, and
$(U_{A}\otimes U_{B}\otimes U_{C}\otimes U_{D})|\Psi\rangle=\sum_{j=0}^{R-1}\lambda_{j}|jjjj\rangle\ ,$ (142)
with $\sum_{j=0}^{R-1}|\lambda_{j}|^{2}=1$. This is precisely the generalized GHZ form. ∎
Because the local QFT cannot have a separable state for the spatial partition, the generalization of the qudit-state proof already implies that the local QFT for the spatial partition does not allow a GHZ-type state if the qudit state can approach the QFT. The GHZ-type state can only explore the mode-wise entanglement, such as in Lifshitz’s vacuum state [2]. We will demonstrate the computability via third-order negativity in CFT2 in the next section.
6 Application in CFT
In this section, we consider the third-order negativity in CFT2. Take a Cauchy slice and choose three ordered points $z_{1}$, $z_{2}$, and $z_{3}$. They divide the slice into three adjacent intervals:
$A=(-\infty,z_{1}]\cup[z_{3},+\infty)\ ,\quad B=[z_{1},z_{2}]\ ,\quad C=[z_{2},z_{3}]\ .$ (143)
We now study the third-order negativity associated with the bipartition $B|C$.
The replica construction of the third-order negativity (16) is represented by a three-point function of twist operators,
$\operatorname{Tr}(\rho_{BC}^{\Gamma})^{3}=\big\langle\sigma_{g_{1}}(z_{1},\bar{z}_{1})\,\sigma_{g_{2}}(z_{2},\bar{z}_{2})\,\sigma_{g_{3}}(z_{3},\bar{z}_{3})\big\rangle\ ,$ (144)
where the permutation elements $g_{j}\in S_{3}$ are determined by the gluing conditions in the replica manifold. For the third-order negativity, each insertion is associated with a $3$-cycle, so the conformal dimensions of the three twist operators are identical,
$\Delta_{1}=\Delta_{2}=\Delta_{3}=\frac{c}{12}\Big(3-\frac{1}{3}\Big)\ ,$ (145)
where $c$ is the central charge. The three-point function of primaries is fixed by conformal symmetry up to the OPE coefficient [18, 28],
$\big\langle\sigma_{1}(z_{1},\bar{z}_{1})\,\sigma_{2}(z_{2},\bar{z}_{2})\,\sigma_{3}(z_{3},\bar{z}_{3})\big\rangle=\frac{C_{123}}{|z_{12}/\epsilon|^{\Delta_{1}+\Delta_{2}-\Delta_{3}}|z_{23}/\epsilon|^{\Delta_{2}+\Delta_{3}-\Delta_{1}}|z_{31}/\epsilon|^{\Delta_{3}+\Delta_{1}-\Delta_{2}}}\ ,$ (146)
where $z_{jk}:=z_{j}-z_{k}$, and $\epsilon$ is the UV cutoff.
In a large-$c$ CFT, the OPE coefficient of the heavy twist operators can be obtained from the AdS/CFT correspondence [29, 22, 46]. When $\Delta_{1}$, $\Delta_{2}$, and $\Delta_{3}$ satisfy the triangle inequalities, which is beyond the CFT computable region, the holographic method provides a consistent result for the OPE coefficient [10]
$\begin{split}\ln C_{123}=&\,\frac{1}{2}\Delta_{1}\ln\bigg(\frac{(\Delta_{1}+\Delta_{2}-\Delta_{3})(\Delta_{1}+\Delta_{3}-\Delta_{2})}{\Delta_{2}+\Delta_{3}-\Delta_{1}}\bigg)+\text{(2\ perm.)}\\ &\,+\frac{1}{2}\big({\textstyle\sum}_{j}\Delta_{j}\big)\big(\ln({\textstyle\sum}_{j}\Delta_{j})-\ln 4\big)-{\textstyle\sum}_{j}\Delta_{j}\ln\Delta_{j}\ .\end{split}$ (147)
Substituting Eq. (145) into Eq. (147) gives
$\ln C_{123}=\frac{c}{3}\ln\frac{3}{4}\ .$ (148)
Combining Eqs. (146), (145), and (148), we arrive at
$\ln\operatorname{Tr}(\rho_{BC}^{\Gamma})^{3}=-\frac{c}{9}\ln\bigg(\frac{|z_{1}-z_{2}|^{2}|z_{2}-z_{3}|^{2}|z_{3}-z_{1}|^{2}}{\epsilon^{6}}\bigg)+\frac{c}{3}\ln\frac{3}{4}\ .$ (149)
This gives the universal large-$c$ result for the third-order negativity of two adjacent intervals in the vacuum state. Thus, multipartite measures are computable quantities in CFT rather than merely formal definitions.
7 Discussion and Conclusion
This paper studied multipartite entanglement through the interplay between separability criteria and entanglement measures. We showed that the third-order negativity $I_{5}$ (7) satisfies
$I_{5}=1$ (150)
if and only if a tripartite pure state is a fully product state. Its convex-roof extension [15] then provides a necessary and sufficient criterion for full separability of tripartite mixed states. This criterion naturally applies to four-qubit pure states and yields four $\phi$-type” tripartite measures. Together with the six two-tangles [3, 24, 49], four three-tangles [12], and four quadripartite measures, these quantities form a complete set of eighteen independent entanglement measures, matching the number of LU-invariant degrees of freedom of a four-qubit pure state. For three-qubit pure states, the upper bound of $\phi_{ABC}$ and the lower bound of $I_{5}$ are shown to be saturated by the GHZ- and W-states, respectively, thereby interpolating between separable states and distinct classes of entangled states. We further proved that the four-tangle [40, 48] is the only entanglement measure that can remain nonzero while all other seventeen vanish, uniquely characterizing generalized GHZ-type four-qubit states. Extending beyond qubits, we developed a replica-based construction of separability criteria for general finite-dimensional multipartite systems, showing that a $\mathtt{q}$-partite pure state is LU-equivalent to a generalized GHZ-state if and only if all of its ($\mathtt{q}-1$)-partite reduced density matrices are fully separable. This establishes a direct link between multipartite separability and global entanglement structure. Finally, we discussed applications to QFT by computing the third-order negativity in CFT2, highlighting the connection between multipartite entanglement measures and field-theoretic observables.
There are several promising directions for future investigation: Although the convex-roof constructions used in this work are conceptually natural, they are generally difficult to evaluate in practice. Developing analytic formulae or efficient numerical methods for entanglement measures such as $\phi_{ABC}$ [32] would significantly enhance the applicability of the separability criteria.
Following Table 1, one might expect that for a five-qubit pure state, there are $\binom{5}{2}=10$ bipartite, $2\times\binom{5}{3}=20$ tripartite, and $4\times\binom{5}{4}=20$ quadripartite entanglement measures. However, this already exceeds the number of independent parameters of a five-qubit pure state [8]:
10+20+20>47.10+20+20>47\ . (151)
A similar mismatch occurs for six qubits,
$\binom{6}{2}+2\times\binom{6}{3}+4\times\binom{6}{4}>2^{6+1}-2-3\times 6\ .$ (152)
This count indicates that the four-qubit system is the last case in which the naive hierarchical organization remains compatible with parameter independence. For five qubits and beyond, multipartite measures must satisfy nontrivial constraints and hidden relations. Identifying these relations and extracting an independent minimal set of measures is an important open problem.
In this work, almost all four-qubit entanglement measures are expressed in terms of reduced density matrices, with the notable exception of the hyperdeterminant. Our results indicate that the hyperdeterminant is nonvanishing only for specific classes of genuinely multipartite entangled states. It would therefore be valuable to express the hyperdeterminant directly in terms of reduced density matrices, as this could provide deeper insight into the structure of four-qubit entanglement.
More generally, many-body systems exhibit qualitatively different types of entanglement. It is therefore important to develop measures—such as $\phi_{ABC}$—that can detect how a separable state transitions into distinct classes of entangled states. Since any valid entanglement measure must vanish on separable states, the study of separability provides a natural guiding principle for constructing such measures. Moreover, our framework applies to finite-dimensional Hilbert spaces. It is directly accessible to current experimental platforms, particularly in the four-qubit pure-state setting, which represents the simplest realization of tripartite mixed states.
Finally, it is important to explore how these ideas extend to QFT. A common approach is to approximate QFT by qudit systems in the large local-dimension limit. However, our results suggest a limitation of this correspondence: generalized GHZ-type qudit states are separable under all reduced density matrices, whereas local QFT does not admit strictly separable states for spatial subregions [47]. This indicates that not all entanglement structures realized in qudit systems are compatible with local QFT. It would be interesting to characterize which classes of states remain consistent with the intrinsic entanglement structure of QFT. This question is closely related to the fact that, while qudit systems admit a standard partial trace, local QFT lacks a straightforward tensor-factor decomposition for spatial regions [47]. Consequently, qudit systems may more naturally capture mode-wise entanglement, as exemplified by the vacuum structure of Lifshitz theories [2], rather than spatial entanglement in local QFT.
Acknowledgments
We want to express our gratitude to Chong-Sun Chu, Ruihua Fan, Jiayu Ran, Peng Tan, and Huangjun Zhu for their helpful discussion. CTM thanks Nan-Peng Ma for his encouragement.
Appendix A Lower Bound of $I_{5}$
Proposition 8.
For any three-qubit pure state, $I_{5}\geq 2/9$, and the lower bound is saturated if and only if the state is LU-equivalent to the $W$ state $(|100\rangle+|010\rangle+|001\rangle)\sqrt{3}$.
Proof.
Any three-qubit pure state is LU-equivalent to the following canonical form [1, 4]
$|\psi\rangle=\lambda_{0}|000\rangle+\lambda_{1}e^{i\varphi}|100\rangle+\lambda_{2}|101\rangle+\lambda_{3}|110\rangle+\lambda_{4}|111\rangle\ ,$ (153)
with $\lambda_{i}\geq 0$, $0\leq\varphi<\pi$, and $\sum_{j=0}^{4}\lambda_{j}^{2}=1$. Let
$a=\lambda_{0}^{2}\ ,\quad b=\lambda_{1}^{2}\ ,\quad c=\lambda_{2}^{2}\ ,\quad d=\lambda_{3}^{2}\ ,\quad e=\lambda_{4}^{2}\ ,$ (154)
so that $a,b,c,d,e\geq 0$ and $a+b+c+d+e=1$. Substituting Eq. (153) directly into the definition of $I_{5}$ (7) gives
$I_{5}=$ $\,a^{3}+b^{3}+c^{3}+d^{3}+e^{3}$ (155)
$+3(a^{2}b+ab^{2}+b^{2}c+bc^{2}+b^{2}d+bd^{2}+c^{2}e+ce^{2}+d^{2}e+de^{2})$
$+3(abc+abd+acd+bcd+bce+bde+cde)$
$+6\sqrt{bcde}\,(b+c+d+e)\cos\varphi\ .$
Since we have $b+c+d+e=1-a\geq 0$, the last term in Eq. (155) is minimized, at fixed $a,b,c,d,e$, by taking $\cos\varphi=-1$. Now define
[…]
so that
$0\leq y\leq\frac{s}{2}\ ,\quad e=1-a-b-s\ .$ (157)
Using Eq. (156), one finds after simplification
$I_{5}\geq 1+3(a+b)^{2}+3bs-3a-3b+3(2a-1)y^{2}-6(1-a)\sqrt{be}\,y\ .$ (158)
Thus, the problem is reduced to minimizing
$\mathcal{I}_{5}\equiv 1+3(a+b)^{2}+3bs-3a-3b+3(2a-1)y^{2}-6(1-a)\sqrt{be}\,y$ (159)
in the compact domain:
$a,b,s,y\geq 0\ ,\quad a+b+s\leq 1\ ,\quad 0\leq y\leq\frac{s}{2}\ .$ (160)
We first exclude interior critical points. Assuming
[…]
we then get
$\frac{\partial\mathcal{I}_{5}}{\partial s}=3b+\frac{3(1-a)b}{\sqrt{be}}y>0\ .$ (162)
Therefore, there is no interior critical point. Hence, the global minimum must lie on the boundary of the domain (160). We now analyze the boundary faces.
[…]
Then Eq. (158) reduces to
$\mathcal{I}_{5}=1+3(a+b)^{2}-3(a+b)+3bs\geq 3(a+b)^{2}-3(a+b)+1\geq\frac{1}{4}>\frac{2}{9}\ .$ (163)
[…]
Then Eq. (158) becomes
$\mathcal{I}_{5}=1+3b^{2}+3bs-3b-3y^{2}-6\sqrt{be}\,y\ .$ (164)
Since the coefficients of both $y^{2}$ and $y$ are negative, the minimum occurs at the largest allowed value $y=s/2$. Hence, we get
$\mathcal{I}_{5}=1+3b^{2}+3bs-3b-\frac{3}{4}s^{2}-3s\sqrt{be}\ .$ (165)
Using
$\sqrt{be}\leq\frac{b+e}{2}=\frac{1-s}{2}\ ,$ (166)
we then get
$\mathcal{I}_{5}\geq 1+3b^{2}+3bs-3b-\frac{3}{4}s^{2}-\frac{3}{2}s(1-s)=\frac{1}{4}+3\left(b-\frac{1-s}{2}\right)^{2}\geq\frac{1}{4}>\frac{2}{9}\ .$ (167)
[…]
Then Eq. (158) gives
$\mathcal{I}_{5}=1-3a+3a^{2}+3(2a-1)y^{2}\ .$ (168)
If $a\geq 1/2$, then $2a-1\geq 0$, so the minimum occurs at $y=0$, and hence we get
$\mathcal{I}_{5}\geq 1-3a+3a^{2}\geq\frac{1}{4}\ .$ (169)
If $a\leq 1/2$, then $2a-1\leq 0$, so the minimum occurs at the largest allowed value $y=s/2$. Since also $s\leq 1-a$, the minimum is attained at s=1as=1-a. Therefore, we obtain
$\mathcal{I}_{5}\geq 1-3a+3a^{2}+\frac{3}{4}(2a-1)(1-a)^{2}=\frac{6a^{3}-3a^{2}+1}{4}\ .$ (170)
On $0\leq a\leq 1/2$, the RHS is minimized at $a=1/3$, giving
$\mathcal{I}_{5}\geq\frac{2}{9}\ .$ (171)
Equality holds if and only if
$a=\frac{1}{3}\ ,\quad s=\frac{2}{3}\ ,\quad y=\frac{1}{3}\ ,$ (172)
which implies
$a=c=d=\frac{1}{3}\ ,\quad b=e=0\ .$ (173)
[…]
Then Eq. (158) reduces to
$\mathcal{I}_{5}=1-3as+3(2a-1)y^{2}\ .$ (174)
If $a\geq 1/2$, then the minimum is at $y=0$, and since $s\leq 1-a$,
$\mathcal{I}_{5}\geq 1-3a(1-a)\geq\frac{1}{4}\ .$ (175)
If $a\leq 1/2$, then the minimum is at $y=s/2$, and again $s\leq 1-a$, so
$\mathcal{I}_{5}\geq 1-3a(1-a)+\frac{3}{4}(2a-1)(1-a)^{2}=\frac{6a^{3}-3a^{2}+1}{4}\geq\frac{2}{9}\ .$ (176)
Equality holds if and only if
$a=\frac{1}{3}\ ,\quad s=\frac{2}{3}\ ,\quad y=\frac{1}{3}\ ,$ (177)
which again implies Eq. (173).
[…]
Then necessarily $y=0$, so this is contained in case (i).
Combining all boundary cases, we conclude that
$I_{5}\geq\frac{2}{9}\ .$ (178)
Moreover, equality can occur only in cases (iii) and (iv), namely Eq. (173). This is precisely the W-state in the canonical form. Therefore, the lower bound is saturated if and only if the state is LU-equivalent to the W-state. ∎
Appendix B Upper Bound of $\phi_{ABC}$
Proposition 9.
For any three-qubit pure state, $\phi_{ABC}\leq 99/2$, and the upper bound is saturated if and only if the state is LU-equivalent to the GHZ-state $(|000\rangle+|111\rangle)/\sqrt{2}$.
Proof.
Using the conventions in Eqs. (153) and (154), a direct substitution into the definition of $\phi_{ABC}$ (9) gives
$\frac{\phi_{ABC}}{18}=11a(1-a)-11ab+8be-3a(c+d)+(8-4a)cd-4(4-a)\sqrt{bcde}\cos\varphi\ .$ (179)
Since 4a>04-a>0, the RHS is maximized, at fixed $a,b,c,d,e$, by taking $\cos\varphi=-1$. Hence, we get
$\frac{\phi_{ABC}}{18}\leq 11a(1-a)-11ab+8be-3a(c+d)+(8-4a)cd+4(4-a)\sqrt{bcde}\ .$ (180)
Next fix $a,b,e$ and $s:=c+d$. The dependence on $c,d$ in the RHS of Eq. (180) is
$3as+(8-4a)cd+4(4-a)\sqrt{becd}\ .$ (181)
Because $0\leq a\leq 1$, both coefficients (84a)(8-4a) and $4(4-a)\sqrt{be}$ are positive, so this is increasing in $cd$. Since $cd\leq s^{2}/4$ with the equality held at $c=d=s/2$, the maximum is attained when $c=d=s/2$. Therefore, it suffices to maximize
$\Phi(a,b,s)\equiv 11a(1-a)-11ab+8be-3as+(2-a)s^{2}+2(4-a)s\sqrt{be}\ ,$ (182)
where now
[…]
We now analyze the boundary faces of the simplex.
[…]
Then
$\Phi(0,b,s)=2s^{2}+8be+8s\sqrt{be}\ ,\quad b+s+e=1\ .$ (184)
Using $be\leq(1-s)^{2}/4$ and $\sqrt{be}\leq(1-s)/2$, we get
$\Phi(0,b,s)\leq 2s^{2}+2(1-s)^{2}+4s(1-s)=2\ .$ (185)
[…]
Then $a+b+s=1$ and
$\Phi(a,b,s)=11a(1-a)-11ab-3as+(2-a)s^{2}=2s^{2}+8as-as^{2}\ .$ (186)
For fixed $s$, this is linear in $a$, with positive coefficient s(8s)s(8-s) on $[0,1]$, so the maximum occurs at a=1sa=1-s, i.e. $b=0$. Hence, we get
$\Phi\leq s^{3}-7s^{2}+8s\leq\frac{68}{27}\ .$ (187)
[…]
Then
$\Phi(a,0,s)=11a(1-a)-3as+(2-a)s^{2}\ .$ (188)
For fixed $a$, this is convex in $s$, since $2-a\geq 1>0$. Hence, the maximum on $0\leq s\leq 1-a$ occurs at an endpoint:
$\Phi(a,0,0)$ $=11a(1-a)\leq\frac{11}{4}\ ,$ (189)
$\Phi(a,0,1-a)$ $=(1-a)(a^{2}+5a+2)\leq\frac{68}{27}\ .$ (190)
Thus, we obtain
$\max_{b=0}\Phi=\frac{11}{4}\ .$ (191)
[…]
Then
$\Phi(a,b,0)=11a(1-a)-11ab+8b(1-a-b)\ .$ (192)
For fixed $a$, this is concave in $b$, and its maximum is:
$\max_{b}\Phi(a,b,0)=\begin{cases}\dfrac{(8+3a)^{2}}{32}\ ,&0\leq a\leq\dfrac{8}{19}\ ,\\[6.0pt] 11a(1-a)\ ,&\dfrac{8}{19}\leq a\leq 1\ .\end{cases}$ (193)
Therefore, we obtain
$\max_{s=0}\Phi=\frac{11}{4}\ ,$ (194)
attained at $a=1/2$, $b=0$. Hence, every boundary face satisfies
$\Phi\leq\frac{11}{4}\ ,$ (195)
and equality can only occur at
$a=\frac{1}{2}\ ,\quad b=0\ ,\quad s=0\ ,\quad e=\frac{1}{2}\ .$ (196)
It remains to exclude interior critical points. For fixed $a,s$, one computes
$\frac{\partial^{2}\Phi}{\partial b^{2}}=-16-\frac{(4-a)s(1-a-s)^{2}}{2[b(1-a-s-b)]^{3/2}}<0\qquad(0<b<1-a-s)\ .$ (197)
Thus, for fixed $(a,s)$, the function is strictly concave in $b$, so any interior critical point is unique in the $b$-direction. Now assume:
a>0,b>0,s>0,e>0.a>0\ $\quad$ b>0\ $\quad$ s>0\ $\quad$ e>0\ . (198)
Let $r:=\sqrt{b/e}>0$, so that
$b=\frac{(1-a-s)r^{2}}{1+r^{2}}\ ,$ (199)
Rewrite $\Phi(a,b,s)$ in terms of $a,s,r$, which we denote as $\tilde{\Phi}(a,r,s)$. Then the stationarity equations $\partial_{a}\tilde{\Phi}=\partial_{r}\tilde{\Phi}=\partial_{s}\tilde{\Phi}=0$ become polynomial equations in $(a,r,s)$:
$\partial_{a}\tilde{\Phi}=0\ \Rightarrow\$$(1-r)^{2}(1+r^{2})s^{2}+\big((5+6a)r^{2}+22a-11\big)$
$\qquad-(8r^{4}+(4a-10)r^{3}+21r^{2}+(4a-10)r-3)s=0\ ,$ (200)
$\partial_{r}\tilde{\Phi}=0\ \Rightarrow\$$\big((4-a)r^{4}-8r^{3}+8r-(4-a)\big)s+(3a+8)r^{3}+19ar-8r=0\ ,$ (201)
$\partial_{s}\tilde{\Phi}=0\ \Rightarrow\$$\big(4(1-r)^{4}-2a(1-r)^{2}(1+r^{2})\big)s+8r(1-r)^{2}$
$\qquad+2a^{2}r(1+r^{2})+a(8r^{4}-10r^{3}+21r^{2}-10r-3)=0\ .$ (202)
Eliminating $s$ from Eqs. (201) and (202) yields
$2a^{2}r(r+1)+a(8r^{3}-16r^{2}+17r-3)-32r^{3}+76r^{2}-56r+12=0\ .$ (203)
Eliminating $s$ from Eqs. (200) and (201), and then eliminating $a$ between the resulting equation and Eq. (203), one obtains the resultant
$(r-1)^{4}(r+1)(2r-1)(4r-7)(4r^{2}-4r+3)(5r^{3}-r^{2}+21r-17)^{2}=0\ .$ (204)
Hence, the only positive candidates are:
$r=\frac{1}{2}\ ,\quad r=1\ ,\quad r=\frac{7}{4}\ ,\quad r=r_{*}\ ,$ (205)
where $r_{*}$ is the unique positive root of
$5r^{3}-r^{2}+21r-17=0\ .$ (206)
Substituting these candidates back into the stationarity equations (200)-(202) gives:
  • $r=1/2$ implies $(a,s)=(1/3,2/3)$, thus 1as=01-a-s=0, i.e. a boundary point;
  • $r=1$ forces $a=0$, again a boundary point;
  • $r=7/4$ yields no admissible point in the simplex;
  • $r=r_{*}$ yields no real solution with $a,b,s,e>0$.
Therefore, there is no interior critical point.
Since the domain is compact, the global maximum of $\Phi$ must lie on the boundary. By Eq. (195), it follows that $\Phi\leq\frac{11}{4}$, thus we conclude
$\phi_{ABC}\leq\frac{99}{2}\ .$ (207)
Equality holds precisely at
$a=e=\frac{1}{2}\ ,\quad b=c=d=0\ ,$ (208)
namely for the GHZ-state. ∎